### v1

# SAMM: Sharded Automated Market Makers

Abstract
Automated Market Makers (AMMs) are a cornerstone of decentralized finance (DeFi) blockchain-based platforms. They are smart contracts, enabling the direct exchange of virtual tokens by maintaining liquidity pools. Traders exchange tokens with the contract, paying a fee; liquidity comes from liquidity providers, paid by those fees. But despite growing demand, the performance of AMMs is limited. State-of-the-art blockchain platforms allow for parallel execution of transactions. However, we show that AMMs do not enjoy these gains, since their operations are not commutative so transactions using them must be serialized.
We present SAMM, an AMM comprising multiple independent shards. All shards are smart contracts operating in the same chain, but they allow for parallel execution as each is independent. The challenge is that trading in a standard AMM is cheaper if its liquidity pool is larger. Therefore, we show that simply using multiple smaller AMMs results in traders splitting each trade among all AMMs, which worsens performance. SAMM addresses this issue with a novel design of the trading fees. Traders are incentivized to use only a single smallest shard. We show that all Subgame-Perfect Nash Equilibria (SPNE) fit the desired behavior: Liquidity providers balance the liquidity among all pools, so the system converges to the state where trades are evenly distributed.
Evaluation in the Sui blockchain shows that SAMM’s throughput is over fivefold that of traditional AMMs, approaching the system’s limit. SAMM is a directly deployable open-source smart contract, allowing trading at scale for individuals and DeFi applications.
1 Introduction
[…]
Figure 1: Trade transaction latency as a function of demand with $n$ SAMM shards.
In this work, we address the throughput limitation of AMMs by using multiple AMM instances called shards. We model the system (§3) as a set of AMM shards and rational users of two kinds. Traders purchase tokens, they use the available AMMs and pay fees as required, aiming to minimize their expenses. Liquidity providers deposit tokens into AMMs and earn fees based on their contribution.
The shards are AMMs based on the standard Constant Product Market Maker (CPMM) contract (§4). Roughly, the contract maintains the product of the two tokens constant after each trade. Thus, purchasing a larger amount of a token increases its unit cost, an effect called slippage.
We present SAMM (§5), an AMM protocol that uses multiple shards, all of which are AMM smart contracts operating on the same blockchain. Ideally, the shards should be balanced, i.e., have equal liquidity (deposited amounts), and traders should randomly select a shard to complete each trade.
The model gives rise to a game (§6) played among the users. In each step, either a liquidity provider adds liquidity to a subset of the shards, or a trader executes a trade using a subset of the shards. We assume myopic liquidity provider behavior, reducing the analysis to a Stackelberg game where the liquidity provider adds liquidity to maximize her revenue from a subsequent trade. We observe that naively using a set of independent CPMMs results in all trades being split among all CPMMs, increasing system overhead without improving throughput. To overcome this, rather than using a set fee ratio (as in all previous work we are aware of), we set a range of possible fees. Within this range, SAMM uses a trading fee function that encourages traders to use the smallest shard.
Our analysis (§7) shows that, indeed, in all best responses, traders use one of the smallest pools. This, in turn, implies that not filling the smallest pool is not the best response for a liquidity provider. We provide specific strategies for traders and liquidity providers that form a subgame perfect equilibrium. We also show that, once the system reaches the balanced state, it will stay in that state.
To evaluate SAMM (§8), we implemented the protocol and deployed it to a local test network of the Sui blockchain platform [10]. SAMM achieves over a fivefold throughput increase compared to a standard single-contract AMM. Figure 1 shows that with more shards, SAMM achieves higher throughput (X axis) with lower trimmed-mean latency (Y axis). Error bars show additional experiments, X marking failure due to overload. This increase is limited by the serial elements of Sui’s transaction processing, following Amdahl’s Law. Finally, we confirm the theoretical analysis by simulating trades from real data and observe that (1) traders follow the desired behavior and (2) SAMM significantly improves the liquidity providers’ revenue with a minor increase in the traders’ costs due to enhanced throughput that allows for more trades.
In summary, our contributions are: (1) Identification of the performance challenges due to “hot” contracts, (2) generalization of the trading fee function of AMMs, (3) SAMM: sharded AMM contract with a novel trading fee function to incentivize the desired behavior, (4) game-theoretic analysis showing Subgame-Perfect Nash Equilibrium (SPNE), (5) evaluation in Sui, demonstrating a fivefold increase in throughput (up to the blockchain’s limit), and (6) simulation with real trade data confirming the theoretical analysis.
These results hint at an upcoming challenge (§9) in smart contract platform design: minimizing the serial elements of transaction processing. But SAMM can already be employed to scale AMM performance both for direct usage and as part of DeFi smart contracts.
2 Related Work
The introduction of the constant product market maker (CPMM) model by Uniswap v1 [52] set a new standard for AMMs, employing a liquidity pool and an algorithm designed to keep the product of the token balances constant. This approach enabled asset exchanges without relying on traditional order books. Subsequent iterations, Uniswap v2 [4] and v3 [3], further developed the CPMM model by improving the price oracle mechanism and the returns for liquidity providers, respectively. As a result, the CPMM algorithm has become a benchmark, with many AMMs adopting similar trading mechanisms [34, 44, 15].
Academic research has primarily concentrated on theoretical models, utility optimization, and security issues surrounding AMMs. Angeris et al.[5] expanded the understanding of AMMs by delving into constant function market makers (CFMMs), demonstrating their utility as decentralized price oracles and broadening the CPMM model’s application. Following this, research has increasingly focused on trading utility maximization [6], advanced arbitrage techniques [53, 26, 7, 46], improving liquidity providers’ returns [32, 23], ensuring transaction privacy [14], eliminating Miner Extractable Value (MEV) for fair trades [13, 48, 12], and examining the synergy between blockchain-based AMMs and prediction market mechanisms [40]. To the best of our knowledge, previous work did not address AMM throughput scaling.
Like other smart contracts, AMMs’ throughput is limited by the blockchain’s constraints. AMM contracts can be deployed on so-called layer-2 solutions[2] (e.g., ZkSwap [27] and QuickSwap [37]), but this merely creates a separate environment for AMM contracts, with scaling issues persisting within this realm. Thus, the efficiency of AMM contracts largely depends on improvements in the underlying blockchain.
The first generation of blockchains, starting from Bitcoin [33], suffered from throughput limitations due to their consensus protocols. However, a body of work overcame this limitation using a variety of protocols [18, 51, 49, 43, 16] and data structures  [29, 38]. With consensus constraints out of the way, the serial execution of blockchain transactions became the bottleneck.
Several works propose blockchain sharding [50, 25, 47], i.e., dividing the blockchain into smaller, interconnected chains (shards) allowing parallelism. In a similar vein, layer-2 protocols [24, 8, 11] outsource computation to a secondary protocol secured by the main blockchain [35]. However, both sharding and layer-2 solutions only parallelize independent contracts. So an AMM does not benefit from blockchain sharding or layer-2 solutions, as it must be located in a single chain and processed sequentially. Note that in SAMM we use multiple AMM contracts, which can be run on a single-shard blockchain, or in separate shards of a sharded blockchain.
An alternative approach identifies read and write set conflicts and parallelizes non-conflicting smart contract transaction execution [42, 17, 31, 36, 21, 10]. However, AMM transactions must be processed sequentially and do not benefit from this approach either.
3 Model
We abstract away the blockchain details for our analysis and model the system as a set of participants (§3.1) interacting directly with AMMs (§3.2), exchanging tokens. The system progresses in discrete steps (§3.3) and there exists an external market used by arbitrageurs (§3.4) to arbitrage in our system. The model uses a generic AMM, which we will later instantiate based on previous work and with SAMM.
3.1 Participants
There are two types of participants, liquidity providers and traders. Each liquidity provider holds some token A and token B. They aim to increase their holdings. Traders are either AB or BA, based on their goals. Each BA trader occasionally wishes to get a certain amount of token A, and vice versa for AB traders. They have sufficiently many tokens of the opposite type to complete their trade, but they aim to minimize the cost of obtaining the desired tokens. Both liquidity providers and traders can send and receive tokens to and from the smart contracts. We ignore the gas fee of smart contract executions for simplicity.
3.2 Automated Market Makers
The system also includes Automated Market Makers, automatons that facilitate depositing and trading of tokens. Each AMM maintains some deposited amounts $R^{A}$ of token A and $R^{B}$ of token B. We call these tokens liquidities. Consequently, we call the AMM a liquidity pool.
Within an AMM pool, there are three primary operations: liquidity addition, liquidity removal, and trade.
  • Liquidity Addition: the liquidity provider deposits $I^{A}$ token A and $I^{B}$ token B to the contract.
  • Liquidity Removal: the liquidity provider withdraws $O^{A}$ token A and $O^{B}$ token B from the contract.
  • Trade: The trader sends $I^{A}$ token A (or, alternatively, $I^{B}$ token B) to the contract and gets $O^{B}$ token B (resp., $O^{A}$ token A) from the contract.
In a trade operation, the required input amount for a trader to receive a specific output amount of another token depends on both the output amount and the AMM’s current state. We define the gross amount of a BA trader as the required input amount of token B to get $O^{A}$ token A in the AMM. We denote it by $G(R^{A},R^{B},O^{A})$ (resp., $G(R^{B},R^{A},O^{B})$ for AB traders). Within the gross amount, traders pay a so-called trading fee which contributes to the liquidity providers’ revenue.
Liquidity providers supply liquidities by depositing their tokens in the contract. Once contributing to the pool, a provider receives tokens from trading fees, hence earning revenue. Liquidity providers can later withdraw their tokens from the pool.
We follow prior studies [32, 23] and assume that the trading fee is directly paid to the liquidity providers. Although many practical AMMs (e.g., Uniswap v2[4]) reinvest the trading fees into the liquidity pool, allowing liquidity providers to withdraw more tokens than they deposited as utilities, the trading fee’s impact is negligible relative to the deposited amount. The average ratio of the output amount to the deposited amount is minimal (less than $0.036\%$ as we find in Appendix B), and automated market makers (AMMs) typically charge a low trading fee relative to the gross amount (specifically, $0.3\%$ in Uniswap), so the trading fee’s impact is negligible relative to the amount of deposited tokens. Therefore, our model remains applicable even if trading fees were to be reinvested, given their negligible size.
3.3 System State and Progress
In the system, there are $n$ independent AMM liquidity pools $\textit{pool}_{1},\textit{pool}_{2},\cdots\textit{pool}_{n}$. The system progresses with discrete steps $k=0,1,2,\cdots$ and is orchestrated by a scheduler. In each step, the scheduler randomly selects a participant and this participant executes transactions. The probability of choosing a liquidity provider is $P_{lp}\geq 0$, while the probability of choosing a trader is $P_{t}\geq 0$, with $P_{lp}+P_{t}=1$.
The scheduler assigns the liquidity provider $l^{A}$ token A and $l^{B}$ token B, where $(l^{A},l^{B})$ follows a random distribution $D_{lp}$.
The trader is either a BA trader aiming to obtain token A or an AB trader aiming to obtain token B. The probability of drawing an AB trader (resp., BA trader) is $P_{t}^{\textit{AB}}\geq 0$ (resp., $P_{t}^{\textit{BA}}\geq 0$), with $P_{t}^{\textit{AB}}+P_{t}^{\textit{BA}}=P_{t}$. To avoid repetition, we only show the case of BA traders, the expressions for AB traders are symmetric. The system assigns the BA trader an amount $b^{\textit{BA}}$ token A to obtain following a random distribution $D^{\textit{BA}}$.
3.4 External Market and Arbitrageurs
Following previous work [23, 32], we assume there is an external market providing the price of token A and token B, $p^{A}$ and $p^{B}$, respectively. These prices do not change due to trades; they serve as objective prices for token A and token B.
When someone can get a lower price of tokens in the AMM than the external market, then he can buy tokens from the AMM and sell them in the external market to make profits or vice versa if the price is higher in the AMM; this is arbitrage. Previous work [32, 23, 12, 13] assume active and rational arbitrageurs who can use the external market and always make arbitrages to maximize their utility. We follow the assumption [32] that there are immediate arbitrages without trading fees when the token price in the AMM is different from the external market.
3.5 Our Goal
The throughput of a single AMM contract is limited due to the underlying blockchain. Our goal is to design a set of AMM contracts to improve the overall throughput of the system, despite the individual rational behavior of all participants.
4 Preliminaries: CPMM
As a baseline for SAMM, we consider the Constant Product Market Maker (CPMM, e.g. [3, 34, 44, 15]), the prominent AMM in real-world applications. It uses a share-based solution to manage liquidity addition and removal operations (§4.1) and keeps the product of the deposited amount of two tokens constant in trade operations (§4.2). Liquidity providers earn revenue from the trading fees of traders (§4.3).
4.1 Liquidity Addition and Removal
Most CPMMs (e.g., [52, 4, 34, 37]) use a fungible share token to manage liquidity addition and removal operations. These tokens represent a liquidity provider’s share in the pool.
When liquidity providers add tokens to an AMM, they receive share tokens which signify their portion of the pool. To recall, $R^{A}$ and $R^{B}$ denote the amounts of token A and token B already deposited in the pool. Similarly, $I^{A}$ and $I^{B}$ represent the quantities of token A and token B that the liquidity provider contributes through a liquidity addition operation. Let $R^{S}$ represent the total amount of all share tokens distributed before the operation. The amount of share tokens acquired by the liquidity provider in this operation, $O^{S}$, is given by:
$O^{S}=R^{S}\times\min\left\{\frac{I^{A}}{R^{A}},\frac{I^{B}}{R^{B}}\right\}\,\,.$
This term $\min\left\{\frac{I^{A}}{R^{A}},\frac{I^{B}}{R^{B}}\right\}$ signifies the ratio of the input token to the deposited token. The $\min$ function serves to ensure that the ownership accurately reflects the liquidity provider’s contribution relative to the scarcer asset. It prevents situations where a liquidity provider inputs a large amount of a certain token to unfairly obtain a larger share of tokens in the pool.
Liquidity providers have the option to withdraw tokens from the pool with the liquidity removal operation, which takes input share tokens and outputs token A and token B. Let $I^{S}$ represent the amount of input share tokens, and $R^{S}$ denote the total amount of all share tokens referred to the pool before the execution. The amounts of token A and token B withdrawn are $O^{A}$ and $O^{B}$:
$O^{A}=\frac{I^{S}}{R^{S}}\times R^{A},O^{B}=\frac{I^{S}}{R^{S}}\times R^{B}\,\,.$
Note that $R^{S}$ is not the amount of share tokens deposited in the AMM pool, but the total amount of share tokens owned by all liquidity providers.
4.2 CPMM Trades
Recall that in a trade operation, a trader sends $I^{A}$ token A (resp., $I^{B}$ token B) and gets $O^{B}$ token B (resp., $O^{A}$ token A). A trade is thus defined by a tuple $(I^{A},O^{A},I^{B},O^{B})$, where all values are non-negative, $I^{A},O^{A},I^{B},O^{B}\geq 0$. The trade is either Token $A$ for Token $B$ or Token $B$ for Token $A$, i.e., $I^{A}=O^{B}=0\text{ or }I^{B}=O^{A}=0$.
After the trade, the amount of deposited tokens is updated to $R^{A}+I^{A}-O^{A}$ and $R^{B}+I^{B}-O^{B}$, respectively. Ignoring fees, the CPMM chooses the outputs ($O^{A}$ and $O^{B}$) by setting an invariant called the trading function $\Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})$ [5] which is the product of the amount of token A and token B after the trade, i.e.,
$\Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})\mathrel{\mathop{:}}=\\ (R^{A}+I^{A}-O^{A})\times(R^{B}+I^{B}-O^{B})\,\,.$
Note that $\Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},0,0,0,0)$ is the product of the amount of token A and token B before the trade.
A trade $(I^{A},O^{A},I^{B},O^{B})$ is legal if the trading function remains constant, i.e.,
$\Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})=\Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},0,0,0,0)\,\,.$ (1)
It indicates that the product of the amounts of token A and token B after the trade is the same as the product of the amounts before the trade, hence then names Constant Product Market Maker, i.e.,
$(R^{A}+I^{A}-O^{A})\times(R^{B}+I^{B}-O^{B})=R^{A}\times R^{B}\,\,.$ (2)
If the trader gets $O^{A}$ token A (resp., $O^{B}$ token B), according to Equation 2, she pays
$I^{B}=\frac{R^{A}\times R^{B}}{R^{A}-O^{A}}-R^{B}=\frac{R^{B}\times O^{A}}{R^{A}-O^{A}}\,\,,$ (3)
and similarly for an AB trader.
Note that we ignored fees in the above equations. We call the payment without fees (Equation 3) net amount, and denote by
$\textit{net}(R^{A},R^{B},O^{A})=\frac{R^{B}\times O^{A}}{R^{A}-O^{A}}\,\,.$ (4)
Denote the amount of token B that the trader needs to pay to get a single token A by $p^{AB}=\frac{I^{B}}{O^{A}}$. From the above equation, $p^{AB}=\frac{R^{B}}{R^{A}-O^{A}}$. This value increases as the output amount of token A, $O^{A}$, increases, which is the Slippage of the trade. When the output amount of token B, $O^{B}$, approaches zero, the token price is not influenced by the slippage. We call it the reported price of token A relative to token B and denote it by
$p_{\textit{reported}}^{AB}\mathrel{\mathop{:}}=\lim_{O^{A}\to 0}\frac{R^{B}}{R^{A}-O^{A}}=\frac{R^{B}}{R^{A}}\,\,.$ (5)
When the reported price of an AMM is different from the price in the external market without trading fees, i.e. $p_{\textit{reported}}^{AB}\neq\frac{p^{A}}{p^{B}}$, there is an arbitrage opportunity for arbitrageurs to make profits. Therefore, due to the arbitrageurs, the reported price of the AMM is always equal to the price in the external market without trading fees [32]. That is:
$\frac{p^{A}}{p^{B}}=p_{\textit{reported}}^{AB}=\frac{R^{B}}{R^{A}}\,\,.$ (6)
Since trading fees are not added to the pool (as defined in Section 3.2), the product of $R^{A}$ and $R^{B}$ remains constant after each trade (Equation 2). Then, arbitrageurs keep the ratio of $R^{A}$ and $R^{B}$ equal to $\frac{p^{A}}{p^{B}}$ (Equation 6). Therefore, $R^{A}$ and $R^{B}$ remain the same after the trade and arbitrage.
4.3 CPMM Trading fee
AMMs charge a trading fee for each trade operation, which is paid by the trader. These trading fees form the revenue of liquidity providers. In CPMMs, the trading fee is a constant fraction ${1-\gamma\in[0,1]}$ of input tokens [5]. This is achieved by selecting the trading function $\Phi_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})$ as [5]
$\Phi_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})\mathrel{\mathop{:}}=\\ (R^{A}+\gamma I^{A}-O^{A})\times(R^{B}+\gamma I^{B}-O^{B})\,\,.$
Since a trade $(I^{A},O^{A},I^{B},O^{B})$ is legal if the trading function remains constant (Equation 1), to get $O^{A}$ token A, the trader pays the gross amount
$G_{\textit{CPMM}}(R^{A},R^{B},O^{A})$ $=\frac{1}{\gamma}\left(\frac{R^{A}\times R^{B}}{R^{A}-O^{A}}-R^{B}\right)$
$=\frac{1}{\gamma}\left(\frac{R^{B}\times O^{A}}{R^{A}-O^{A}}\right)\,\,.$ (7)
Compared to the net amount (Equation 3), the trader pays additional tokens to complete the trade; this is the trading fee. In the CPMM case, it is
$\frac{1-\gamma}{\gamma}\left(\frac{R^{B}\times O^{A}}{R^{A}-O^{A}}\right)\,\,.$ (8)
In the prominent Uniswap v2 [4], the ratio is ${1-\gamma=0.003}$.
5 SAMM: Sharded AMM
We introduce SAMM, the first sharded Automated Market Maker. We explain the driving rationale and derive two desired properties of the contract (§5.1). We present the SAMM trading fee function (§5.2) and find parameters to fullfill both properties (§5.3).
5.1 SAMM Structure and Properties
We enable parallel processing of AMM operations by deploying multiple AMM shards. There should not be any data dependencies among the shards and no global elements.
Most operations in an AMM are trade operations ($99.5\%$ based on publicly available blockchain records, see Appendix B). Our main goal is therefore that traders distribute the workload evenly among the shards. This boils down to two properties.
5.1.1 c-non-splitting property
First, each trade should use a single shard only, and not split the trade into smaller ones on multiple shards. That is, the cost of a single transaction should be less than the combined cost of multiple, split transactions. However, this principle faces a significant challenge as highlighted by Equation 3: when the output amount is not small enough in comparison to the shard’s reserve amount, the resulting slippage could incentivize traders to split their transactions to mitigate this slippage. Consequently, we refine our requirement to ensure that this principle is adhered to only when the output amount is relatively small compared to the deposited amount, specifically when the ratio is below a predefined constant, $c$. Indeed, the prominent pairs in Ethereum’s Uniswap v2 data support this approach, with $99\%$ of trades having a ratio of output amount to deposited amount below $0.0052$ (see Appendix B for more details). We call this the $c$-Non-Splitting property.
{restatable}
[$c$-Non-Spliting]proppropertyconcavity Let $m\geq 2$. Given a set of output amount by $\{O^{A}_{j}|1\leq j\leq m,O^{A}_{j}>0\}$, denote by $\tilde{O}^{A}$ the sum of the amounts in the set, $\tilde{O}^{A}=\sum_{j=1}^{m}O^{A}_{j}$. For the constant $10<c<$ 1 and the deposited amount of tokens $R^{A}$ and $R^{B}$, if $\frac{\tilde{O}^{A}}{R^{A}}\leq c$ , then the cost of trading $\tilde{O}^{A}$ token A is less than the sum of the cost of trading $O^{A}_{j}$ token A for $1\leq j\leq m$, i.e.,
$G_{\textit{SAMM}}(R^{A},R^{B},\tilde{O}^{A})<\sum_{j=1}^{m}G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{j})$
5.1.2 c-smaller-better property
Our second goal is to maintain balanced volumes in all shards. This is crucial because the volume directly affects the slippage. When there are stark differences in pool sizes, with some being much smaller than others, the slippage in trading within these smaller pools is significantly greater than in larger ones. This discrepancy can result in transactions clustering in the larger pools rather than spread evenly, reducing parallelism. We address this by incentivizing liquidity providers to allocate their tokens to the shards with lower volumes. Intuitively, this ensures that smaller pools receive more frequent fees from traders when the volumes of pools are not balanced, which incentivizes liquidity providers to deposit tokens in these smaller pools. Therefore, the system would converge to the state where all pools have balanced volumes, and traders then randomly select pools for trading. Similar to the $c$-Non-Splitting property, large transactions suffer from high slippage, which leads to a strong advantage of larger pools. So here too, we refine this requirement to scenarios where the ratio of the traded amount and the deposited amount is below a threshold $c$. We call this the $c$-smaller-better property.
{restatable}
[$c$-smaller-better]proppropertysmallerpoolsmallercost Given an output amount ${O^{A}>0}$, for any two pools with deposited token amounts $(R^{A}_{i},R^{B}_{i})$ and $(R^{A}_{j},R^{B}_{j})$, respectively, and $R^{A}_{i}<R^{A}_{j}$. For the constant $10<c<$ 1, if $\frac{O^{A}}{R^{A}_{j}}<\frac{O^{A}}{R^{A}_{i}}\leq c$ and $\frac{R^{A}_{i}}{R^{B}_{i}}=\frac{R^{A}_{j}}{R^{B}_{j}}$, then the cost of trading $O^{A}$ token A in the smaller pool is less than that in the larger pool, i.e.,
$G_{\textit{SAMM}}(R^{A}_{i},R^{B}_{i},O^{A})<G_{\textit{SAMM}}(R^{A}_{j},R^{B}_{j},O^{A})\,\,.$
5.1.3 $c$ value
If the above properties are satisfied for a particular $c$ but a trade occurs with a larger ratio of output amount to the deposited amount, traders may split their transactions or tend to larger pools to minimize their cost. Therefore, $c$ should be as large as possible to ensure such occurrences are rare.
Note
Finally, we observe that the traditional CPMM cost function, $G_{\textit{CPMM}}(R^{A},R^{B},O^{A})$, does not satisfy either property (Appendix C).
5.2 Trading Fee Design
To satisfy properties 5.1.1 and 5.1.2, we first generalize the trading fee function to provide flexibility in the incentive design (§5.2.1). Then, we propose a specific trading fee function (§5.2.2).
5.2.1 Generic Trading Fee Function
The gross amount of a trade comprises the trading fee and the net amount, with the latter being determined by the CPMM curve. To maintain the foundational characteristics of AMMs, such as reported price, we do not modify the CPMM curve. Instead, we generalize the trading fee function beyond simply taking a ratio of the net amount as in previous work (e.g., [5, 23, 6]).
Denote the trading fee function of the AMM by $\textit{tf}(R^{A},R^{B},O^{A})$, which takes the deposited amount of token A, $R^{A}$, and token B, $R^{B}$, in the pool and the output amount of token A, $O^{A}$, and outputs the amount of token B the trader needs to pay as the trading fee. Then, the gross amount of getting $O^{A}$ token A is:
$G(R^{A},R^{B},O^{A})=\textit{tf}(R^{A},R^{B},O^{A})+\textit{net}(R^{A},R^{B},O^{A})\,\,.$ (9)
Recall the trading fee of a CPMM is given in Equation 8.
5.2.2 Bounded-Ratio Trading Fee Function
In order to achieve the desired properties, we need flexibility for the trading fee function design. A monomial function is sufficient to achieve most of our goals. The function takes the variables available on a trade, $R^{A},R^{B},O^{A}$. It is parameterized by four values, $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$:
$\textit{tf}(R^{A},R^{B},O^{A};\beta_{1},\beta_{2},\beta_{3},\beta_{4})\mathrel{\mathop{:}}=\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}\,\,.$
While the monomial function offers a straightforward approach to calculating trading fees, its lack of bounds poses a challenge. Without limits, the trading fee might become excessively high, deterring traders, or too low, diminishing the revenue for liquidity providers. To address this, there is a need for adjustable boundaries similar to setting a single fixed ratio in previous work. The limits allow for the fine-tuning of the trading fee’s absolute value. To navigate these concerns, we introduce the bounded-ratio polynomial function based on the monomial, which introduces $r_{\min}$ and $r_{\max}$ as parameters to control the trading fee’s range and $\beta_{5}$ as a parameter to adjust the trading fee’s base value:
$\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A};\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5})\mathrel{\mathop{:}}=\\ \frac{R^{B}}{R^{A}}O^{A}\times\max\{r_{\min},\min\{r_{\max},\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}\}\}$ (10)
The ratio $\frac{R^{A}}{R^{B}}$ represents the market price of token A relative to token B (see Section 4.2). The product $\frac{R^{B}}{R^{A}}O^{A}$ is thus the trader’s net payment in terms of token B without slippage. Then $\max\{r_{\min},\min\{r_{\max},\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{B})^{\beta_{4}}+\beta_{5}\}\}$ acts as the constrained ratio of the trading fee to that net payment. We omit the parameters $\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5}$ in the rest of this paper for brevity.
5.3 Parameter Selection
Now we turn to the selection of parameters for the SAMM trading fee function. First, we identify the necessary conditions under which the bounded-ratio polynomial trading fee function aligns with the $c$-smaller-better property.
{restatable}
propostheoremnecessary Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, then the following conditions are necessary for $c$-smaller-better to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
  1. 1.
    $\beta_{3}=0$,
  2. 2.
    $\beta_{2}+\beta_{4}=0$,
  3. 3.
    $\beta_{1}<0$,
  4. 4.
    $0<\beta_{4}\leq 1$,
  5. 5.
    $r_{\min}<\beta_{5}\leq r_{\max}$, and
  6. 6.
    $\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq c^{\beta_{4}}$.
We require the polynomial value to be between $r_{\min}$ and $r_{\max}$. Then, we need to make sure that the derivative of the gross amount on the size of the pool is non-negative to ensure the $c$-smaller-better property. Required items come directly from these two restrictions. We defer the proof to Appendix D.
From the above theorem, we require that $\beta_{1}<0,\beta_{2}+\beta_{4}=0,\beta_{3}=0,0<\beta_{4}\leq 1,r_{\min}<\beta_{5}\leq r_{\max}$. Next, we identify additional conditions that are sufficient for both properties to hold.
{restatable}
thmtheoremsufficient Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, if $\beta_{1}<0,\beta_{2}+\beta_{4}=0,\beta_{3}=0,0<\beta_{4}\leq 1,r_{\min}<\beta_{5}\leq r_{\max}$ and $\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq c^{\beta_{4}}$, then following items are sufficient for the $c$-Non-Splitting and $c$-smaller-better properties to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
  1. 1.
    $\beta_{1}\beta_{4}(\beta_{4}+1)c^{\beta_{4}-1}(1-c)^{3}\leq-2$
  2. 2.
    $-\beta_{1}\beta_{4}\geq\frac{c^{1-\beta_{4}}}{(1-c)^{2}}$
The $c$-smaller better property is satisfied when the derivative of the gross amount is positive. It is sufficient to ensure the $c$-Non-Splitting property when the gross amount is concave to the output amount, which is ensured by a negative second derivative over the output amount. The proof is in Appendix E.
By setting $\beta_{2}=-1,\beta_{3}=0,\beta_{4}=1,\beta_{5}=r_{\max}$, and choosing $\beta_{1}<-1$, the fee function satisfies the above requirements, leaving just three parameters:
$\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\frac{R^{B}}{R^{A}}\times O^{A}\times\max\left\{r_{\min},\beta_{1}\times\frac{O^{A}}{R^{A}}+r_{\max}\right\}\,\,.$
By setting $\beta_{4}=1$ and $\beta_{5}=r_{\max}$ in the sufficient condition of the $c$-Non-Splitting and $c$-smaller-better properties (Theorem 5.3), the sufficient condition becomes: {restatable}cortheoremspecialpara For any $\beta_{1}<-1$ and $c$ satisfying
$c\leq\min\left\{1-(-\beta_{1})^{-\frac{1}{3}},\frac{r_{\max}-r_{\min}}{-\beta_{1}}\right\},$
the SAMM cost function $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$ satisfies the $c$-Non-Splitting and $c$-smaller-better properties.
For instance, the parameters $\beta_{1}=-1.05$, $r_{\max}=0.012$, $r_{\min}=0.001$, and $c=0.0104$ meet the specified criteria. According to historical records (Appendix B), over $99\%$ of Uniswap v2 transactions have a ratio below $0.0052$, suggesting that if these pools are split into two shards, $99\%$ of transactions would fall within our targeted range. Specifically, in the pools with the highest trading volumes, USDC-ETH and USDT-ETH, the ratio remains below $0.00128$, which can manage eight shards. Increasing the number of shards could be achieved by adjusting the value of $c$ by increasing $r_{\max}$ and $\beta_{1}$.
6 Game-theoretic Analysis
The model gives rise to a game (§6.1) played among traders and liquidity providers. Then, we find a specific property that AMM algorithms should exhibit to enhance throughput (§6.2).
6.1 Game Model
Our model gives rise to a sequential game with discrete steps $k=0,1,2,\cdots$. The game is parameterized by the number $n$ of pools and by trading fee functions tf of the AMMs. We denote it by $\Gamma_{n}(\textit{tf})$.
6.1.1 System State
In $\Gamma_{n}(\textit{tf})$, the state of each pool $\textit{pool}_{i}$ consists of the amount of deposited token A, the amount of deposited token B and the amount of share tokens, $R^{A}_{i},R^{B}_{i},R^{S}_{i}$ respectively. Recall that share tokens are not deposited in the pool but are held by liquidity providers and $R^{S}_{i}$ is the total amount of its related share tokens held by liquidity providers. We denote the state of all AMM contracts in step $k$ by ${\textbf{R}({k})=\left(\left(R_{1}^{A}({k}),R_{1}^{B}({k}),R_{1}^{S}({k})\right),\cdots\left(R_{n}^{A}({k}),R_{n}^{B}({k}),R_{n}^{S}({k})\right)\right)}$.
6.1.2 Liquidity Provider Actions
The liquidity provider decides the amount of tokens she deposits in each AMM pool. We denote the amount of token A and token B depositted in $\textit{pool}_{i}$ by $l^{A}_{i},l^{B}_{i}\geq 0$, respectively. Recall that the scheduler assign the liquidity provider $l^{A}$ token A and $l^{B}$ token B. The total amount of tokens deposited should not exceed the amount she holds:
$\forall 1\leq i\leq n,l^{A}_{i},l^{B}_{i}\geq 0,\sum_{i=1}^{n}l^{A}_{i}\leq l^{A},\sum_{i=1}^{n}l^{B}_{i}\leq l^{B}\,\,.$
The action of a liquidity provider is thus the vector $a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)$. When the liquidity provider takes this action with the system state $\textbf{R}({k})$, the liquidity provide receives $O^{S}_{i}$ share token from $\textit{pool}_{i}$, where
$O^{S}_{i}=R_{i}^{S}({k})\times\min\left\{\frac{l^{A}_{i}}{R_{i}^{A}({k})},\frac{l^{B}_{i}}{R_{i}^{B}({k})}\right\}\,\,.$
Since $\frac{R^{A}_{i}}{R^{B}_{i}}=\frac{p^{B}}{p^{A}}$ (Equation 6), when $\frac{l^{A}_{i}}{l^{B}_{i}}=\frac{p^{B}}{p^{A}}$, solely increasing $l^{A}_{i}$ or $l^{B}_{i}$ would not increase the share token the liquidity provider receives, which means more payment without more revenue. Therefore, we only consider actions where $\frac{l^{A}_{i}}{l^{B}_{i}}=\frac{p^{B}}{p^{A}}$ and require $\frac{l^{A}}{l^{B}}=\frac{p^{B}}{p^{A}}$. The action space of a liquidity provider is denoted by $\mathcal{A}_{\textit{lp}}(l^{A},l^{B})$; it is the set of all feasible actions:
$\mathcal{A}_{lp}(l^{A},l^{B})=\\ \left\{a_{lp}\left|\forall 1\leq i\leq n,l^{A}_{i}=\frac{p^{B}}{p^{A}}l^{B}_{i}\geq 0,\sum_{i=1}^{n}l^{A}_{i}\leq l^{A},\sum_{i=1}^{n}l^{B}_{i}\leq l^{B}\right.\right\}\,\,.$
We denote the updated state of AMM pools from the previous state R and the action of a liquidity provider $a_{lp}$ by $\textbf{R}+a_{lp}$. Then for $\textbf{R}^{\prime}=\left(\left(R^{A^{\prime}}_{1},R^{B^{\prime}}_{1},R^{S^{\prime}}_{1}\right),\cdots\left(R^{A^{\prime}}_{n},R^{B^{\prime}}_{n},R^{S^{\prime}}_{n}\right)\right)=\textbf{R}+a_{lp}$, we have
$R^{A^{\prime}}_{i}=R^{A}_{i}+l^{A}_{i}\,\,,$
$R^{B^{\prime}}_{i}=R^{B}_{i}+l^{B}_{i}\,\,,$
$R^{S^{\prime}}_{i}=R^{S}_{i}+O^{S}_{i}=\left(1+\frac{l^{A}_{i}}{R^{A}_{i}}\right)R^{S}_{i}\,\,.$
After the liquidity addition operation, the reported price of $\textit{pool}_{i}$ (Equation 5) becomes $p_{\textit{reported}}^{AB}=\frac{R^{A^{\prime}}_{i}}{R^{B^{\prime}}_{i}}=\frac{R^{A}_{i}+l^{A}_{i}}{R^{B}_{i}+l^{B}_{i}}=\frac{R^{A}_{i}}{R^{B}_{i}}=\frac{p^{B}}{p^{A}}$, which is the price in the external market. Therefore, there is no arbitrage opportunity for the arbitrageurs. Then the update of state in step $k+1$ is
$\textbf{R}({k+1})=\textbf{R}({k})+a_{lp}\,\,.$
6.1.3 Trader Actions
The action of a BA trader determines the amount of token A she acquires from each AMM pool. Denote by $b^{BA}_{i}\geq 0$ the amount of token A she acquires in $\textit{pool}_{i}$. Recall that the scheduler assigns the BA trader $b^{\textit{BA}}$ token A to acquire in total. The action of a BA trader is thus the vector $a^{\textit{BA}}=\left(b^{\textit{BA}}_{1},\cdots,b^{\textit{BA}}_{n}\right)$. The action space of a BA trader is denoted by $\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})$, which is the set of all feasible actions:
$\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})=\left\{a^{\textit{BA}}\left|\forall 1\leq i\leq n,b^{\textit{BA}}\geq 0,\sum_{i=1}^{n}b^{\textit{BA}}_{i}=b^{\textit{BA}}\right.\right\}\,\,.$ (11)
Recall that after the trade operation and the arbitrage, $R^{A}_{i}$ and $R^{B}_{i}$ remain unchanged (Section 4.2). Consequently, the state of the liquidity pools remains unchanged in the subsequent step:
$\forall 1\leq i\leq n,R_{i}^{A}({k+1})=R_{i}^{A}({k}),R_{i}^{B}({k+1})=R_{i}^{B}({k})\,\,.$ (12)
6.1.4 Utility and Strategies
For traders and liquidity providers, we first discuss their revenue and then define their strategies and utility, respectively. We assume that traders and liquidity providers consider the value of tokens the same as the prices of the external market, namely $p^{A}$ and $p^{B}$.
Consider a BA trader whose goal is to acquire $b^{\textit{BA}}$ units of token A. This trader needs to pay the gross amount and may derive some fixed reward from getting these tokens. We consider her revenue only as the inverse of the gross amount in terms of token B times the value of each token B:
$U^{\textit{BA}}(\textbf{R},a^{\textit{BA}})=$ $-p^{B}\times\sum_{i}G(R^{A}_{i},R^{B}_{i},b^{\textit{BA}}_{i})$
$=$ $-p^{B}\times\sum_{i}G(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}}_{i})\,\,.$ (13)
For the trader aiming to get $b^{\textit{BA}}$ token A, the strategy of the trader $\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ takes R, $b^{\textit{BA}}$ and an action $a^{\textit{BA}}$ as input, then outputs the probability of taking action $a^{\textit{BA}}$. The total probability of all feasible actions should be 1,
$\sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})}\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=1\,\,.$ (14)
The utility of the trader over the strategy is a function of the system state R, the assigned requirement $b^{\textit{BA}}$, and the strategy of the trader $\pi^{BA}$. It is the expected revenue under the distribution of actions,
$U^{BA}(\textbf{R},b^{\textit{BA}},\pi^{BA})=\\ \sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})}\left(\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})\times U^{BA}(\textbf{R},a^{\textit{BA}})\right)\,\,.$ (15)
The revenue of liquidity providers comes from the trading fees paid by traders. We consider the myopic setting (as in e.g. [39, 19]) where the liquidity provider would measure her utility in the next step as her long-term revenue.
Denote the revenue of a liquidity provider with her action $a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)$, the action of the BA trader in the next step $a^{\textit{BA}}=\left(b^{\textit{BA}}_{1},\cdots,b^{\textit{BA}}_{n}\right)$ and the system state R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$, by the function $U_{lp}(\textbf{R},a_{lp},a^{BA})$. In the next step, $\textit{pool}_{i}$ receives a trading fee of $\textit{tf}(R^{A}_{i}+l^{A}_{i},R^{B}_{i}+l^{B}_{i},b^{\textit{BA}}_{i})$. The liquidity provider receives a fraction of that fee proportional to her fraction of share tokens out of all shares in the pool. Therefore, the revenue function is:
$U_{lp}(\textbf{R},\,a_{lp},a^{BA})$
$=$ $p^{B}\times\sum_{i=1}^{n}\left\{\textit{tf}(R^{A}_{i}+l^{A}_{i},R^{B}_{i}+l^{B}_{i},b^{\textit{BA}}_{i})\times\frac{\frac{l^{A}_{i}}{R^{A}_{i}}R^{S}_{i}}{\left(1+\frac{l^{A}_{i}}{R^{A}_{i}}\right)R^{S}_{i}}\right\}$
$=$ $p^{B}\times\sum_{i=1}^{n}\left\{\textit{tf}(R^{A}_{i}+l^{A}_{i},R^{B}_{i}+l^{B}_{i},b^{\textit{BA}}_{i})\times\frac{l^{A}_{i}}{l^{A}_{i}+R^{A}_{i}}\right\}\,\,.$ (16)
For the liquidity provider with $l^{A}$ token A and $l^{B}$ token B, the strategy of the liquidity provider $\pi_{lp}(\textbf{R},l^{A},l^{B})$ takes R, $l^{A}$, $l^{B}$ and an action $a_{lp}$ as input, and outputs the probability of taking action $a_{lp}$. The total probability of all feasible actions should be 1,
$\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})=1\,\,.$ (17)
The utility of the liquidity provider over strategies is a function of the system state R, the amount of tokens she is assigned $l^{A}$, $l^{B}$, and the strategy of the liquidity provider and traders, $\pi_{lp},\pi^{BA},\pi^{AB}$; denote it by $U_{lp}(\textbf{R},l^{A},l^{B},\pi_{lp},\pi^{BA},\pi^{AB})$. Before calculating this, we show the revenue given the action of the liquidity provider and the strategies of traders, denoted by $U_{lp}(\textbf{R},l^{A},l^{B},a_{lp},\pi^{BA},\pi^{AB})$. It takes the system state R, the action of the liquidity provider $a_{lp}$, the strategies of traders $\pi^{BA}$ and $\pi^{AB}$ as input, then outputs the expected utility over the strategies and distributions of traders. The strategy of traders is affected by the state after the liquidity provider’s action, namely $\textbf{R}+a_{lp}$. Denote by $E_{b^{\textit{BA}}\sim D^{BA}}[f(\cdot)]$ the expected value of $f(\cdot)$ with $b^{\textit{BA}}$ is sampled from $D^{BA}$. The revenue is:
$U_{lp}(\textbf{R},l^{A},l^{B},a_{lp},\pi^{BA},\pi^{AB})=\\ P_{t}^{\textit{BA}}\times E_{b^{\textit{BA}}\sim D^{BA}}\left[\sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})}\left(\begin{subarray}{c}\pi^{BA}(\textbf{R}+a_{lp},b^{\textit{BA}},a^{\textit{BA}})\times\\ U_{lp}(\textbf{R},a_{lp},a^{BA})\end{subarray}\right)\right]+\\ P_{t}^{\textit{AB}}\times E_{b^{\textit{AB}}\sim D^{AB}}\left[\sum_{a^{\textit{AB}}\in\mathcal{A}_{\textit{AB}}(b^{\textit{AB}})}\left(\begin{subarray}{c}\pi^{AB}(\textbf{R}+a_{lp},b^{\textit{AB}},a^{\textit{AB}})\times\\ U_{lp}(\textbf{R},a_{lp},a^{AB})\end{subarray}\right)\right]\,\,.$
To simplify the presentation, we assume the liquidity provider is always followed by a BA trader. The expressions for an AB trader are symmetric. Then, the above equation can be simplified as
$U_{lp}(\textbf{R},l^{A},l^{B},a_{lp},\pi^{BA})=\\ E_{b^{\textit{BA}}_{i}\sim D^{BA}}\left[\sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})}\left(\begin{subarray}{c}\pi^{BA}(\textbf{R}+a_{lp},b^{\textit{BA}},a^{\textit{BA}})\times\\ U_{lp}(\textbf{R},a_{lp},a^{BA})\end{subarray}\right)\right]\,\,.$ (18)
Then, the utility function of the liquidity provider over strategies is the expected utility under the distribution of actions:
$U_{lp}(\textbf{R},l^{A},l^{B},\pi_{lp},\pi^{BA},\pi^{AB})=\\ \sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\left(\begin{subarray}{c}\pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\\ \times U_{lp}(\textbf{R},l^{A},l^{B},a_{lp},\pi^{BA},\pi^{AB})\end{subarray}\right)\,\,.$ (19)
6.1.5 Solution Concept
The subgame perfect Nash equilibrium (SPNE) ensures that in sequential games, players cannot gain higher utility by changing strategies at any game step, supported by backward induction [41] where players optimize their utility based on previous actions.
When a trader takes an actions in a given step, her utility is influenced solely by her immediate strategy and the current state of AMMs, as outlined in Equation 6.1.4. Crucially, future actions do not bear on this calculation allowing the trader to directly optimize her utility function, thereby establishing dominant strategies.
In the case of a liquidity provider being chosen in a step, the situation is different. Given their myopic viewpoint, liquidity providers only need to account for the strategy of the trader in the ensuing step. Their actions in subsequent steps do not affect their own utility. Thus the sequential game is reduced to a two-stage Stackelberg game and SPNE to a Stackelberg Equilibrium [45].
To formalize this, we denote the strategies of the liquidity provider, the BA trader, and the AB trader in the SPNE by $\tau_{lp},\tau^{BA}$ and $\tau^{AB}$, respectively. The BA trader would always get the optimal utility in equilibrium, namely $\forall\textbf{R},b^{\textit{BA}}$, we have
$U^{\textit{BA}}(\tau^{BA},\textbf{R},b^{\textit{BA}})=\max_{\pi^{BA}}U^{\textit{BA}}(\pi^{BA},\textbf{R},b^{\textit{BA}})\,\,.$
Note that $\tau^{BA}$ is a best response for the BA trader. The strategy of liquidity provider in equilibrium is just the optimal strategy when traders adopt their best response, namely $\forall\textbf{R},l^{A},l^{B}$, we have
$U_{lp}(\tau_{lp},\textbf{R},\tau^{BA},\tau^{AB},l^{A},l^{B})=\\ \max_{\pi_{lp}}U_{lp}(\pi_{lp},\textbf{R},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$
6.2 Desired Property
Our goal is to improve the throughput by allowing parallelism. specifically, we would like traders to evenly distribute their transactions among all AMM contracts without splitting them. That is, a dominant strategy for the BA trader should be to randomly select an AMM contract to acquire all her needed token A. Denote the action of getting all $b^{\textit{BA}}$ token A in $\textit{pool}_{i}$ by
$a_{i}^{\textit{BA}}(b^{\textit{BA}})=\left(0,\cdots,b^{\textit{BA}}_{i}=b^{\textit{BA}},\cdots,0\right)\,\,.$ (20)
Denote the set of these actions by $\mathcal{A}_{\mathbf{1}}(b^{\textit{BA}})\subset\mathcal{A}_{t}^{\textit{BA}}(b^{\textit{BA}})$:
$\mathcal{A}_{\mathbf{1}}(b^{\textit{BA}})=\left\{a_{i}^{\textit{BA}}(b^{\textit{BA}})\left|1\leq i\leq n\right.\right\}\,\,.$
The strategy that uniformly at random selects an AMM contract to acquire all her needed token A is the perfect parallelism strategy: {restatable}defiperfectparallelismstrategy The perfect parallelism strategy of the BA trader is $\hat{\tau}^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$, where
$\hat{\tau}^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}\frac{1}{n},&\text{if }a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1}}(b^{\textit{BA}})\\ 0,&\text{Otherwise.}\end{cases},$
Our goal is thus to have the perfect parallelism strategy be a dominant strategy:
{restatable}
proppropertyperfectparallelism The perfect parallelism strategy of the BA trader is a dominant strategy, namely
$\forall\pi^{BA}:U^{\textit{BA}}(\pi^{BA},\textbf{R},b^{\textit{BA}})\leq U^{\textit{BA}}(\hat{\tau}^{BA},\textbf{R},b^{\textit{BA}})\,\,.$
Using multiple CPMMs does not satisfy the perfect parallelism property and would be counterproductive: Each trader would split her transactions among all AMM contracts. Thus, although the total number of trades increases due to parallelism, the satisfied trade demand is not higher than a single AMM contract and possibly lower since the total throughput might only increase sublinearly in the number of AMM contracts. Appendix F provides details of this analysis.
7 SAMM Equilibrium
We analyze the behavior of players in the game with SAMM. We first prove the trader randomly selects a pool to trade when the states of pools are balanced (§7.1). Then, we show the system tends to the balanced state since liquidity providers invest their tokens in the smallest pools, reducing the difference in the volume of pools (§7.2). Full proofs are in appendix G.
7.1 Trader Strategy
Consider the case that the system state is R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$. As discussed in Section 5.3, the SAMM gross amount satisfies the $c$-non-splitting property and $c$-smaller-better property for a certain $10<c<$ 1. We assume that the required amount of token A, $b^{\textit{BA}}$, is at most a fraction $c$ of the amount of deposited token A in all pools, i.e.,
$\forall 1\leq i\leq n,b^{\textit{BA}}\leq cR^{A}_{i}\,\,.$
7.1.1 Traders’ optimal action
The $c$-non-splitting property and $c$-smaller-better property give a trader the incentive to randomly select one of the smallest pools to trade all her required tokens. Recall that $a_{i}^{\text{BA}}(b^{\textit{BA}})$ is the action of acquiring all $b^{\textit{BA}}$ token A in $\textit{pool}_{i}$ (Equation 20). We define the set of actions that trade in one of the smallest pools: {restatable}defismallestpoolactionset The Smallest Pool Action Set is the set of actions that acquire all $b^{\textit{BA}}$ token A in one of the smallest pools under state R:
$\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})=\left\{a_{i}^{\textit{BA}}(b^{\textit{BA}})|\forall j,R^{A}_{i}\leq R^{A}_{j}\right\}\,\,.$
The cardinality of $\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$ is the number of smallest pools in R. We denote this by
$n_{\min}(\textbf{R})=\left|\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\right|\,\,.$
When the trader selects one of the actions in the smallest pool action set $\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, she gets the highest revenue:
{restatable}
lemlemmaoptimaltraderaction In $\Gamma_{n}(tf_{\textit{SAMM}})$, a trader wants to get $b^{\textit{BA}}$ token A when the system state is R. Then for the action which obtains all $b^{\textit{BA}}$ token A in one of the smallest pools with index $i^{\ast}$, where $a_{i^{\ast}}^{\text{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, the trader has no less than the revenue of any other actions:
$\forall a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}}),U^{\textit{BA}}(a_{i^{\ast}}^{\text{BA}},\textbf{R})\geq U^{\textit{BA}}(a^{\textit{BA}},\textbf{R})\,\,.$
Proof Sketch.
Due to the $c$-non-splitting property, trading in a single pool is better than trading in multiple pools. Then the revenue of trading in one of the smallest pools is no less than that in any other pool due to the $c$-smaller-better property. ∎
7.1.2 Using smallest pools is a dominant strategy
Lemma 7.1.1 indicates that when multiple AMM pools have the same smallest amount of deposited tokens, acquiring all tokens in any one of them has the highest utility. Since the utility of a trader’s strategy is the linear combination of the utility of actions, it is a dominant strategy for the trader to randomly select one of the smallest pools to acquire all required tokens: {restatable}corlemmadominanttraderstrategy In $\Gamma_{n}(tf_{\textit{SAMM}})$, a dominant strategy of a BA trader is to randomly select one of the smallest pools to acquire all required tokens:
$\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}\frac{1}{n_{\min}(\textbf{R})},&\text{if }a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\\ 0,&\text{Otherwise.}\end{cases}$
If $n_{\min}(\textbf{R})=n$, then all pools have the same amount of deposited tokens, and the trader randomly selects one of the $n$ pools. {restatable}cortheoremperfectparallelism In $\Gamma_{n}(tf_{\textit{SAMM}})$, the system state is R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$. If $\forall i,j,R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, then the perfect parallelism strategy
$\hat{\tau}^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}\frac{1}{n},&\text{if }a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1}}(b^{\textit{BA}})\\ 0,&\text{Otherwise.}\end{cases}$
is a dominant strategy for the BA trader.
7.1.3 All dominant strategies use smallest pools
We have shown that only trading in one of the smallest pools is the dominant strategy for the trader. However, to later determine the best response of liquidity providers, we need to know whether there are other dominant strategies. We show that if the trader has a positive probability of taking the action of splitting a transaction or trading in a pool with not the smallest amount of deposited tokens, then she has strictly lower utility than randomly selecting one of the smallest pools to trade: {restatable}thmtheoremstrictsingle In $\Gamma_{n}(tf_{\textit{SAMM}})$, considering the following dominant strategy of the BA trader which randomly selects one of the smallest pools to acquire all required tokens:
$\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}\frac{1}{n_{\min}(\textbf{R})},&\text{if }a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\\ 0,&\text{Otherwise.}\end{cases},$
then for all strategies $\pi^{BA}$ that have a positive probability of actions not trading in one of the smallest pools, i.e., $\exists a^{\textit{BA}}=\left(b^{\textit{BA}}_{1},\cdots,b^{\textit{BA}}_{i},\cdots,b^{\textit{BA}}_{n}\right)\notin\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R}),\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})>0$ , the utility of the BA trader is strictly lower than with strategy $\tau^{BA}$:
$U^{\textit{BA}}(\tau^{BA},\textbf{R},b^{\textit{BA}})>U^{\textit{BA}}(\pi^{BA},\textbf{R},b^{\textit{BA}})\,\,.$
Proof Sketch.
Since the utility of the BA trader is a linear combination of the utility of actions, we only need to show that the action of not trading in one of the smallest pools has strictly lower revenue than the action of trading in one of the smallest pools, which can be deduced from $c$-smaller-better property and $c$-non-splitting property. ∎
From the above theorem, all the best responses of the trader should only have a positive probability of taking an action that trades in exactly one of the smallest pools: {restatable}corlemmaoptimaltraderstrategyset Considering any best response strategy $\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ of the BA trader, the strategy should only have a positive probability of taking an action that trades in one of the smallest pools:
$\forall a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})\setminus\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}}),\tau^{BA}(\textbf{R},b^{\textit{BA}},a_{i}^{\textit{BA}})=0\,\,.$
In other words, the sum of the probabilities of all actions that trade in one of the smallest pools should be 1:
$\sum_{a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})}\tau^{BA}(\textbf{R},b^{\textit{BA}},a_{i}^{\textit{BA}}(b^{\textit{BA}}))=1\,\,.$ (21)
7.2 Liquidity Provider Strategy and SPNE
Now we turn to the strategies of the liquidity providers and identify the SPNE of the game.
7.2.1 Scaffolding
We want the liquidity provider to fill up smaller pools to keep pools balanced. We call such an action the fillup action, where if the liquidity provider inputs tokens in a pool, then the pool is the smallest pool after this action. We denote the fillup action by $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$: {restatable}defideffillup_action The fillup action of a liquidity provider $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})=\left(\left(\hat{l}^{A}_{1},\hat{l}^{B}_{1}\right),\cdots,\left(\hat{l}^{A}_{n},\hat{l}^{B}_{n}\right)\right)$ is the action satisfying that if the liquidity provider inputs tokens in a pool, then the pool is one of the smallest pools after this action:
$\forall 1\leq i\leq n:\hat{l}^{A}_{i}\geq 0,$
$\sum_{i=1}^{n}\hat{l}^{A}_{i}=l^{A}\,\,;$
$\forall\hat{l}^{A}_{i}>0,\forall j:\hat{l}^{A}_{i}+R^{A}_{i}\leq\hat{l}^{A}_{j}+R^{A}_{j}\,\,.$
We also define the strategy that only takes the fillup action as the fillup strategy: {restatable}defifillupstrategy The fillup strategy of a liquidity provider $\tau^{\textit{fill}}_{lp}(\textbf{R},l^{A},l^{B})$ is the strategy that only takes the fillup action:
$\tau^{\textit{fill}}_{lp}(\textbf{R},l^{A},l^{B},a_{lp})=\begin{cases}1,&\text{if }a_{lp}=\hat{a}_{lp}\\ 0,&\text{Otherwise.}\end{cases}\$
Denote the minimal amount of deposited token A among all pools of status R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$ by
$\rho^{A}(\textbf{R})=\min_{1\leq i\leq n}R^{A}_{i}\,.$
We show that $a_{lp}^{\textit{fill}}$ is unique and has the maximal volume of the smallest pool in the next step.
{restatable}
lemlemmauniquefillup For any action of liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$, if $\rho^{A}(\textbf{R}+a_{lp})\geq\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})$, then $a_{lp}=a_{lp}^{\textit{fill}}$.
Proof Sketch.
If another action results in a higher minimum reserve of token A, then this action must input a larger amount of tokens into each pool compared to the fill-up action, contrary to the assumption that actions have identical total input amounts. ∎
We have shown that traders are incentivized to trade in smaller pools in Section 7.1, which incentivizes liquidity providers to input their tokens in smaller pools. Additionally, if the liquidity provider makes small pools larger, she would get more trading fees, which further incentivizes them to add liquidity to small pools:
{restatable}
lemlemmatradingfee For any two pools $i$ and $j$, if $R^{A}_{i}<R^{A}_{j}$, for any output amount $b^{\textit{BA}}$ of token B, the trading fee of $\textit{pool}_{i}$ is strictly smaller than the trading fee of $\textit{pool}_{j}$:
$tf_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}})<tf_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}})\,\,.$
Proof Sketch.
Due to the $c$-smaller-better property, the gross amount in a larger pool is larger than that in a smaller pool. However, the net amount of a larger pool is smaller than that of a smaller pool (Equation 3). Therefore, the trading fee of a larger pool is larger than that of a smaller pool since the gross amount is the sum of the net amount and the trading fee. ∎
7.2.2 Perfect parallelism under balanced pools
When all pools have identical sizes, the fillup action is to input tokens in all pools evenly, which is the best response of the liquidity provider:
{restatable}
thmtheoremperfectparallelismandlp Denote by $\hat{a}_{lp}=\left(\left(\frac{1}{n}l^{A},\frac{1}{n}l^{B}\right),\cdots\right)$ the action of evenly depositing tokens in all pools. In $\Gamma_{n}(tf_{\textit{SAMM}})$, if for all $i$ and $j$ that the liquidity amounts are the same, $R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, the liquidity provider strategy which only takes action $\hat{a}_{lp}$,
$\tau_{lp}(\textbf{R},l^{A},l^{B},a_{lp})=\begin{cases}1,&\text{if }a_{lp}=\hat{a}_{lp}\\ 0,&\text{Otherwise.}\end{cases}\$
and any best response of the trader constitutes an SPNE.
Proof Sketch.
Given that traders prefer trading in smaller pools while larger pools generate higher trading fees, the liquidity provider should increase their share in the smallest pools and enhance their sizes. This dual objective is optimally achieved by uniformly distributing tokens across all pools. ∎
The above theorem indicates that the liquidity provider keeps the same amount of deposited tokens in the pool after her action. Therefore, the system always works in a state where all pools have the same amount of deposited tokens.
Since randomly choosing one of the smallest pools to trade is a dominant strategy for the trader, the system always works in perfect parallelism:
{restatable}
corcorollaryalwaysstable Denote by $\hat{a}_{lp}=\left(\left(\frac{1}{n}l^{A},\frac{1}{n}l^{B}\right),\cdots\right)$ the action of evenly depositing tokens in all pools. In $\Gamma_{n}(tf_{\textit{SAMM}})$, if for all $i$ and $j$ that the liquidity amounts are the same, $R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, the liquidity provider strategy which only takes action $\hat{a}_{lp}$,
$\tau_{lp}(\textbf{R},l^{A},l^{B},a_{lp})=\begin{cases}1,&\text{if }a_{lp}=\hat{a}_{lp}\\ 0,&\text{Otherwise.}\end{cases}\$
and the BA trader strategy of randomly selecting one of the smallest pools to trade,
$\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}\frac{1}{n_{\min}(\textbf{R})},&\text{if }a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\\ 0,&\text{Otherwise.}\end{cases}$
constitute an SPNE.
7.2.3 Convergence to Balanced Pools
We now show that even if the system reaches an unbalanced state, it converges to the perfect parallelism since the liquidity provider takes the fillup strategy. We can conclude that the fillup strategy is the only best response in all SPNE:
{restatable}
thmtheoremalwaysfillup In $\Gamma_{n}(\textit{tf}_{\textit{SAMM}})$, in all SPNE, the liquidity provider’s best response is the fillup strategy:
$\tau^{\textit{fill}}_{lp}(\textbf{R},l^{A},l^{B},a_{lp})=\begin{cases}1,&\text{if }a_{lp}=a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})\\ 0,&\text{Otherwise.}\end{cases}\,\,.$
Figure 2: An example construction of $a_{lp}^{\prime}$.
Proof Sketch.
Given any action $a_{lp}$ that is not the fillup action, we can construct a new action $a_{lp}^{\prime}$ that is strictly better than $a_{lp}$. By ensuring the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ is larger than that in $\textbf{R}+a_{lp}$, we increase trading fees garnered from each transaction. Moreover, this smallest pool is also the smallest before the action, maximizing the liquidity provider’s share. Consequently, the liquidity provider earns higher revenue under $a_{lp}^{\prime}$ than under $a_{lp}$. Thus, any strategy incorporating an action other than the fillup action is not optimal. Figure 2 illustrates an example construction of $a_{lp}^{\prime}$. ∎
7.2.4 Specific SPNE under deviation
Finally, we find a specific SPNE in $\Gamma_{n}(\textit{tf}_{SAMM})$ when pools are not balanced. In this SPNE, if there are multiple smallest pools, the trader uses the smallest pool in the last step. If there is more than one smallest pool in the last step, the trader selects the one with the smallest index. Denote by $i_{\min}(\textbf{R})$ the index of the pool with the smallest amount of deposited token A in R. When $i^{\ast}=i_{\min}(\textbf{R})$, we have
$\forall j,R^{A}_{i^{\ast}}\leq R^{A}_{j}$
$\forall j,R^{A}_{i^{\ast}}=R^{A}_{j}\Rightarrow i^{\ast}\leq j\,\,.$ (22)
If the pool with index $i_{\min}(\textbf{R})$ is the only pool, the BA trader would only trade in that pool. Then the liquidity provider only taking the fill-up action is the best response strategy: {restatable}thmtheoremfillup In $\Gamma_{n}(\textit{tf}_{SAMM})$, assume that the pool state in step $k$ is $\textbf{R}({k})$, $i^{\ast}=i_{\min}(\textbf{R}({k}))$ is the index of the pool with the smallest amount of deposited token A in $\textbf{R}({k})$. Then the trader strategy is to trade in the smallest pool in the last step if it is the smallest one, or randomly select one of the smallest pools, namely,
$\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}1,&\text{if}R_{i^{\ast}}=\rho^{A}(\textbf{R}({k}),l^{A},l^{B})\text{ and }\\ &a^{\textit{BA}}=a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}})\\ \frac{1}{n_{\min}(\textbf{R})},&\text{if }R_{i^{\ast}}=\rho^{A}(\textbf{R}({k}),l^{A},l^{B})\text{ and }\\ &a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\\ 0,&\text{Otherwise.}\end{cases}$
and the liquidity provider’s fillup strategy:
$\tau^{\textit{fill}}_{lp}(\textbf{R},l^{A},l^{B},a_{lp})=\begin{cases}1,&\text{if }a_{lp}=a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})\\ 0,&\text{Otherwise.}\end{cases},$
are an SPNE in step $k$.
Proof Sketch.
Since the trader always trades in one of the smallest pools, $\tau^{BA}$ is a dominant strategy for her. Therefore, we only need to prove that $\tau_{lp}$ is the best response to $\tau^{BA}$. If the liquidity provider does not take the fill-up action, then the trader trades in the smallest pool with a smaller amount of deposited token A than that under the fill-up action. Since larger pools have a higher trading fee under a fixed trade, the utility of the liquidity provider is higher when she takes the fill-up action. ∎
8 Evaluation
To evaluate the performance of SAMM we use the state-of-the-art Sui blockchain (§8.1). We find that the throughput of a single contract AMM is limited (§8.2), and the throughput can be improved by increasing the number of SAMM shards (§8.3). Our analysis shows that further improvement is possible by increasing the platform’s parallelism (§8.4). Simulation using real trading data shows that traders follow the desired behavior and that SAMM significantly improves the liquidity provider revenue with little sacrifice in trader cost (§8.5)
8.1 Experimental Setup
To evaluate the performance of SAMM, we implement it [^1] Link redacted for anonymous submission. in the Move language [9] in Sui [10], a state-of-the-art blockchain that supports parallel execution. Smart contracts in Sui are independent objects, and Sui executes transactions on different objects in parallel. We deploy a local Sui testnet consisting of $4$ validators, which maintain the consensus of the blockchain. We publish transactions, including transferring tokens, deploying AMM contracts, and executing AMM transactions, via an RPC interface. For all reported results, we use a machine with 2T memory and $256$ CPU cores. Several experiments deploying the Sui testnet on one machine and sending transactions on another machine produce similar results.
We run $100$ trader processes, implemented in Rust, which send transactions to smart contracts through the Sui Rust SDK for the RPC interface. Each trader sends transactions in a random interval following an exponential distribution with expectation $\frac{1}{\lambda}$. Note this experiment is only for performance evaluation, so traders follow the perfect parallelism strategy. We vary the overall frequency of transactions by setting different values of the individual $\lambda$ values. In each test, we set a target throughput. we first warm up the system by sending transactions for $500$ seconds and then measure actual frequencies and latencies for the following $100$ seconds. If the latency is stable within the 100 measurement seconds, we report the mean value. Otherwise, or if more than half of the transactions failed, we consider the experiment failed. The latency is always higher than 1 second due to Sui’s consensus protocol. We consider a latency greater than 3 seconds as a failure.
[…]
8.2 Single-Contract Bottleneck
[…]
8.3 SAMM Evaluation
We implement SAMM by modifying the trading fee mechanism in the OmniSwap contract and deploying a varying number of shards (contracts). When a trader sends a transaction, she randomly selects a SAMM contract.
Figure 1 shows the average latency in different frequencies of transaction demands on different numbers of SAMM contracts. The latency in the case of one SAMM contract is indistinguishable from a single OmniSwap contract. In some instances, more than half of the transactions failed, which is marked as X in the graph. As the number of SAMM contracts increases, the system can process higher demand.
To quantify SAMM’s performance enhancements, we evaluate the throughput with a varying number of shards. We set a latency cap of $3$ seconds. As depicted in Figure 1, the latency escalates rapidly once it surpasses $2$ seconds. Therefore, choosing other latency caps beyond $2$ seconds does not significantly affect the results. The throughput with $n$ shards is thus the highest frequency that produced a latency lower than 3 seconds. Figure 3 shows the throughput increases almost linearly at the beginning and then converges to a bound.
[…]
Figure 3: Maximal throughput as a function of the number of SAMM shards.
8.4 Parallelization in the Underlying Platform
[…]
8.5 Simulation of Trading Data
(a) Relative traders’ costs to the external market.
(b) Relative liquidity providers’ revenue to Uniswap.
Figure 4: SAMM revenue and cost with experimental tps results and trading data.
Figure 5: Distribution of transactions in SAMM
(a) The ratio of transaction splits.
(b) Addition trades due to splits.
Figure 6: Trade Splits in SAMM.
To evaluate SAMM in a real trading environment, we simulate its behavior using a performance profile based on our performance evaluation and workload from publicly available trading data of the Ethereum Uniswap v2 USDC-ETH pool, from Ethereum block $12,000,000$ to $19,500,000$ (from 2021-03-08 to 2024-03-23, about 3 years). We simulate Uniswap v2 and SAMM using $1$ to $32$ shards. We test three different values of $c$, namely $0.003,0.005$, and $0.01$, and choose other parameters through an optimization problem (Appendix I). In particular, there are no arbitrageurs, unlike the theoretical model.
We run each simulation instance as follows. We randomly select a point in Uniswap v2’s history as the starting point. For SAMM with $n$ shards, we evenly distribute the liquidity of the Uniswap v2 contract at that point among the $n$ shards to establish the initial state of SAMM. We then simulate the real trades from that time point onward. Each trade in the real data is simulated as a trading demand with the required amount of tokens, matching the output amount and tokens of the actual trade. To minimize costs, the trader selects the shard offering the lowest price and may split the trade into several smaller trades if this reduces her costs, disregarding gas fees. We assume both the Uniswap and SAMM pools operate at maximal throughput, consistent with the throughput in our performance evaluation. We count trade splits, the number of transactions in different SAMM shards, liquidity provider revenue, and trader costs over 1 second, following a 1-second warm-up period. We repeat each simulation 100 times and calculate averages. Note that $1$ second in our simulation corresponds to several hours in the real world. We get historical prices of ETH to USDC from the Yahoo Finance webpage [^2] https://finance.yahoo.com/quote/ETH-USD/history/ and use it as the external market price. We calculate the revenue for liquidity providers and costs for traders in USDC, converting ETH amounts at the real-time price.
We first analyze the incentives of liquidity providers and traders in SAMM. Figure 4(a) shows the average ratio of traders’ costs in SAMM compared to the costs in the external market, considering different numbers of shards and varying values of $c$. We observe that with larger $c$, the cost for traders increases, as expected. For a fixed $c$, the cost for traders decreases as the number of shards increases, aligning with SAMM’s $c$-smaller-better property, where more shards result in less liquidity in each shard. Additionally, we compare the cost for traders in Uniswap v2, depicted by the black line. In all cases, the cost of SAMM is either smaller ($c=0.003$ with no less than 2 shards) or slightly larger than that of Uniswap v2, differing by less than $1\%$ with $c=0.01$ and $0.3\%$ with $c=0.005$, when there are at least 2 shards. We aggregate trading fees across all shards to evaluate the revenue of liquidity providers. Figure 4(b) shows SAMM consistently generates higher revenue than Uniswap v2, with $c=0.01$ and 7 shards yielding over 15 times the revenue. Initially, revenue in SAMM rises with more shards due to increased throughput but with even more shards it declines as trading fees per trade decrease in smaller shards, as explained in Lemma 7.2.1.
We monitor the distribution of trades across each shard in the SAMM system. Figure 5 illustrates the number of trades executed in each shard compared to the average. Error bars show the range of relative trade numbers in each shard compared to the average. We see where the difference from average is always under $5\%$. Additionally, we analyze the proportion of trades that are split into multiple transactions within SAMM. Figure 6 shows that with $c=0.01$ trade splits are infrequent, occurring in less than $8\%$ of trades, even with 32 shards. Furthermore, the total number of trade splits results in less than a $45\%$ increase in overall trade volume, a relatively minor increment given the five-fold throughput improvement. When the number of shards is larger or $c$ is smaller, the proportion of trade splits is higher.
In summary, our simulation demonstrates that SAMM significantly outperforms Uniswap v2 in terms of liquidity provider revenue while maintaining comparable costs for traders. This increase without major cost hikes is due to SAMM’s enhanced throughput, allowing for more trades and higher total trading fees. The simulation also confirms that our theoretical predictions align with the actual trading behavior. Additionally, the results suggest that SAMM requires a larger $c$ for more shards to prevent trade splits. Although larger $c$ leads to larger trader costs, the cost remains acceptable due to the increased number of shards.
9 Conclusion
We present SAMM, a scalable AMM. The key enabler of SAMM is the design of a trading fee mechanism that incentivizes parallel operations. We analyze trader and liquidity provider behaviors as a game, showing that parallel operations are the best response, and validate by simulation with real trade traces. We implement and deploy SAMM in a local Sui testnet, demonstrating more than 5x throughput improvement, up to the underlying system’s limits. Our results indicate that reducing serial bottlenecks of independent contracts should be a focus of smart-contract platforms to allow for AMM scaling (See Appendix H). Meanwhile, SAMM can be directly deployed to scale AMMs on existing platforms, for direct use and as part of the DeFi eco-system.
Appendix A Growing Demand for AMM
We analyze the demand for AMMs by calculating the number of trades per second on the prominent Uniswap versions 1-3, from its deployment in November 2018 to March 2024. We use data from Dune [^3] https://dune.com/, a comprehensive database for blockchain data. Figure 7 illustrates the exponentially increasing demand, from $0.78$ average trades per second in 2020 to $9.54$ in 2024. On the Ethereum blockchain, most Uniswap transactions are executed through versions 2 and 3, incurring gas costs of $152,809$ and $184,523$ respectively. Given that the average total gas per block is $15,000,000$ and the block interval is $12$ seconds, Ethereum can facilitate up to $8.18$ trades per second for v2 and $6.77$ trades per second for v3. However, in March 2024, the average monthly trades on Uniswap reached $13.9$ per second, nearly doubling Ethereum’s processing capacity. Therefore, most demand of Uniswap is processed through off-chain solutions [2].
[…]
Figure 7: Mothly trades in Uniswap $1$-$3$
Appendix B Uniswap v2 Statistics
We analyze the five Uniswap v2 pools with the highest number of trade transactions from Ethereum block 12,000,000 to 19,500,000 (from 2021-03-08 to 2024-03-23, about 3 years). First, we find that more than $99.5\%$ of the transactions are trade operations. Therefore, we only focus on the throughput of trade operations. Second, we calculate the ratio of output tokens to deposited tokens for each transaction and find that most transactions are small compared with the pool size. Among all trades, the average ratio of output tokens to deposited tokens is less than $0.036\%$, and more than $99\%$ of the trades have a ratio of less than $0.52\%$. Such a phenomenon is consistent in all five pools. Specifically, in the most active pools, USDC-ETH and USDT-ETH, over $99\%$ of trades exhibit a ratio of output to deposited tokens below $0.00128\%$. Figure 8 shows in log scale the $1$ minus the cumulative distribution function of the ratio of output tokens to deposited tokens in all pairs and five selected pairs.
Figure 8: The ratio of output tokens to deposited tokens in Uniswap v2
Appendix C CPMM Does Not Satisfy Either property
Simply deploying multiple CPMM pools does not satisfy our desired properties.
{restatable}
thmthmcpmmlimitation For any value of $10<c<$ 1, the CPMM cost function $G_{\textit{CPMM}}(R^{A},R^{B},O^{A})$ does not satisfy either the $c$-Non-Splitting or the $c$-smaller-better properties.
Proof.
For any $10<c<$ 1, it is sufficient to find a single case where each property does not hold. For the $c$-Non-Splitting property, we consider the cost of getting $\tilde{O}^{A}=cR^{A}$ token A and $O^{A}_{1}=O^{A}_{2}=\frac{c}{2}R^{A}$, the gross amount of getting $\tilde{O}^{A}=cR^{A}$ is larger than getting $O^{A}_{1}$ and $O^{A}_{2}$ respectively:
[…]
$=$ $\frac{1}{\gamma}\left(\frac{R^{B}\times\frac{c}{2}R^{A}}{R^{A}-cR^{A}}\right)+\frac{1}{\gamma}\left(\frac{R^{B}\times\frac{c}{2}R^{A}}{R^{A}-cR^{A}}\right)$
$>$ $\frac{1}{\gamma}\left(\frac{R^{B}\times\frac{c}{2}R^{A}}{R^{A}-\frac{c}{2}R^{A}}\right)+\frac{1}{\gamma}\left(\frac{R^{B}\times\frac{c}{2}R^{A}}{R^{A}-\frac{c}{2}R^{A}}\right)$
[…]
Therefore, the CPMM cost function does not satisfy the $c$-Non-Splitting property.
Next, we turn to the $c$-smaller-better property. Considering two pools with deposited token amounts $(R^{A}_{i},R^{B}_{i})$ and $(R^{A}_{j},R^{B}_{j})$, respectively, where $R^{A}_{i}<R^{A}_{j},\frac{R^{A}_{i}}{R^{B}_{i}}=\frac{R^{A}_{j}}{R^{B}_{j}}$, for any output amount $0<O^{A}\leq cR^{A}_{i}$, consider the gross amount of getting $O^{A}$ token A, we have
[…]
$=$ $\frac{1}{\gamma}\left(\frac{\frac{R^{B}_{i}}{R^{A}_{i}}\times O^{A}}{1-\frac{O^{A}}{R^{A}_{i}}}\right)$
$=$ $\frac{1}{\gamma}\left(\frac{\frac{R^{B}_{j}}{R^{A}_{j}}\times O^{A}}{1-\frac{O^{A}}{R^{A}_{i}}}\right)$
$=$ $\frac{1}{\gamma}\left(\frac{R^{B}_{j}\times O^{A}}{R^{A}_{j}-\frac{R^{A}_{j}}{R^{A}_{i}}O^{A}}\right)$
$>$ $\frac{1}{\gamma}\left(\frac{R^{B}_{j}\times O^{A}}{R^{A}_{j}-O^{A}}\right)$
[…]
Therefore, the CPMM cost function does not satisfy the $c$-smaller-better property as well. ∎
Appendix D Proof of Proposition 5.3
  • Proposition 5.3 (restated). Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, then the following conditions are necessary for $c$-smaller-better to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
    1. 1.
      $\beta_{3}=0$,
    2. 2.
      $\beta_{2}+\beta_{4}=0$,
    3. 3.
      $\beta_{1}<0$,
    4. 4.
      $0<\beta_{4}\leq 1$,
    5. 5.
      $r_{\min}<\beta_{5}\leq r_{\max}$, and
    6. 6.
      $\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq c^{\beta_{4}}$.
Proof.
We would first give some common prefixes for the required items, and then obtain them one by one.
The $c$-smaller-better property considers two pools $i$ and $j$, assuming their reported prices are identical. Denote the inverse of this price by $c^{\textit{AB}}$, so $\frac{R^{A}_{i}}{R^{B}_{i}}=\frac{R^{A}_{j}}{R^{B}_{j}}=\frac{1}{c^{\textit{AB}}}>0$. The property requirement (Equation 5.1.2) becomes
$G_{\textit{SAMM}}(R^{A}_{i},c^{\textit{AB}}R^{A}_{i},O^{A})<G_{\textit{SAMM}}(R^{A}_{j},c^{\textit{AB}}R^{A}_{j},O^{A})\,\,.$
The function $G_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})$ is differentiable, and by assumption $R^{A}_{i}<R^{B}_{i}$, therefore a necessary condition for this inequality to hold is
$\frac{dG_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})}{dR^{A}}\geq 0\,\,.$ (23)
If the polynomial value value is greater than the bound, $\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}>r_{\max}$, then from Equations 49, and 10, the gross amount becomes
$G_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})=c^{\textit{AB}}r_{\max}O^{A}+\frac{c^{\textit{AB}}R^{A}\times O^{A}}{R^{A}-O^{A}}\,\,,$ (24)
so the derivative is
$\frac{dG_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})}{dR^{A}}=-\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-O^{A})^{2}}<0\,\,,$ (25)
contradicting Equation 23. Therefore, for the property to hold we need $\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}\leq r_{\max}$. A similar situation occurs when $\beta_{1}=0$ or $\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}<r_{\min}$, where
$G_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})=\\ c^{\textit{AB}}\max\{r_{\min},\min\{r_{\max},\beta_{5}\}\}O^{A}+\frac{c^{\textit{AB}}R^{A}\times O^{A}}{R^{A}-O^{A}}\,\,,$ (26)
or
$G_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})=c^{\textit{AB}}r_{\min}O^{A}+\frac{c^{\textit{AB}}R^{A}\times O^{A}}{R^{A}-O^{A}}\,\,.$ (27)
Therefore, we also need $\beta_{1}\neq 0$, and $\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}\geq r_{\min}$, to avoid a similar contradiction with Equation 25. Thus, we require $\beta_{1}\neq 0$ and the polynomial to be in the range
$r_{\min}\leq\beta_{1}(R^{A})^{\beta_{2}}(c^{\textit{AB}}R^{A})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}\leq r_{\max}\,\,.$ (28)
Since the output amount $O_{A}$ can be arbitrarily close to zero, for $(O^{A})^{\beta_{4}}$ to be bounded we require
$\beta_{4}\geq 0\,\,.$ (29)
Since $O^{A}$ can be arbitrarily close to zero, $\beta_{1}(R^{A})^{\beta_{2}}(c^{\textit{AB}}R^{A})^{\beta_{3}}(O^{A})^{\beta_{4}}$ can be arbitrarily close to zero, so from Equation 28 we require that
$r_{\min}\leq\beta_{5}\leq r_{\max}\,\,.$ (30)
Here, we start to drive necessary items from the properties. We rewrite Equation 28 as
$r_{\min}\leq\left(c^{\textit{AB}}\right)^{\beta_{3}}\beta_{1}(R^{A})^{\beta_{2}+\beta_{3}+\beta_{4}}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}}+\beta_{5}\leq r_{\max}\,\,.$ (31)
The c-Smaller-Better property should hold for all reported prices, that is, this inequality should hold for all $c^{\textit{AB}}$. But if $\beta_{3}\neq 0$, the first element $\left(c^{\textit{AB}}\right)^{\beta_{3}}$ can be arbitrarily large, and since $\beta_{1}(R^{A})^{\beta_{2}+\beta_{3}+\beta_{4}}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}}\neq 0$, the whole expression $\left(c^{\textit{AB}}\right)^{\beta_{3}}\beta_{1}(R^{A})^{\beta_{2}+\beta_{3}+\beta_{4}}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}}+\beta_{5}$ is unbounded. Therefore, we obtain Item 1:
$\beta_{3}=0\,\,.$ (32)
Similarly, we need to bound the expression $(R^{A})^{\beta_{2}+\beta_{3}+\beta_{4}}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}}$. Since all positive values for $R^{A}$ are possible and all output ratios are bounded $0<\frac{O^{A}}{R^{A}}<c$, to keep the expression bounded for all such values we require $\beta_{2}+\beta_{3}+\beta_{4}=0$, and since we already saw $\beta_{3}=0$, we obtain Item 2:
$\beta_{2}+\beta_{4}=0\,\,.$ (33)
Now, if $\beta_{2}=\beta_{4}=0$, the expression of Equation 28 becomes $\beta_{1}(R^{A})^{\beta_{2}}(c^{\textit{AB}}R^{A})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}=\beta_{1}+\beta_{5}$, so the gross amount is
$G_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})=\\ c^{\textit{AB}}\max\{r_{\min},\min\{r_{\max},\beta_{1}+\beta_{5}\}\}O^{A}+\frac{c^{\textit{AB}}R^{A}\times O^{A}}{R^{A}-O^{A}}\,\,.$ (34)
So, as in Equation 24, the derivative is negative, a contradiction. Therefore, we have $\beta_{4}\neq 0$, and due to Equation 29 we obtain
$\beta_{4}>0,\beta_{2}<0\,\,.$ (35)
Combining the constraints we just found (Equations 32 and 33) into the gross amount expression (Equations 9 and 10) we have
$G_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})=\\ c^{\textit{AB}}O^{A}\times\left(\beta_{1}(R^{A})^{-\beta_{4}}(O^{A})^{\beta_{4}}+\beta_{5}\right)+\frac{c^{\textit{AB}}R^{A}\times O^{A}}{R^{A}-O^{A}}$ (36)
and its derivative is
$\frac{dG_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})}{dR^{A}}=\\ -c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}-\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-O^{A})^{2}}$ (37)
The second element ${\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-O^{A})^{2}}}$ is positive, so to keep the derivative non-negative, the first element $c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}$ must be negative. Since $c^{\textit{AB}}>0$, $\beta_{4}>0$, and $\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}>0$, we obtain Item 3:
$\beta_{1}<0\,\,.$ (38)
Since the derivative is non-negative, from Equation 37 we have
$-c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}\geq\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-O^{A})^{2}}$
Since $O^{A}<c\times R^{A}<R^{A}$, we have $(R^{A}-O^{A})^{2}<(R^{A})^{2}$. Therefore, we have
$-c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}\geq\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A})^{2}}\,\,.$ (39)
By multiplying both sides by $\frac{(R^{A})^{\beta_{4}+1}}{c^{\textit{AB}}(O^{A})^{\beta_{4}+1}}$ which is positive, the above inequality is equal to:
$-\beta_{1}\beta_{4}\geq\left(\frac{O^{A}}{R^{A}}\right)^{1-\beta_{4}}\,\,.$ (40)
Since $\frac{O^{A}}{R^{A}}$ can be arbitrarily close to zero, if $\beta_{4}>1$, the right side of the above inequality can be arbitrarily large, so we require $\beta_{4}<=1$. Combining equation 35, we obtain Item 4:
$0<\beta_{4}\leq 1\,\,.$
From Equation 28, we have
$\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}\geq r_{\min}.$ (41)
And since $\beta_{1}<0$, we have $\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}<0$ so $\beta_{5}>r_{\min}$. This allows us to make the first inequality of Equation 30 strict, which gives us Item 5:
$r_{\min}<\beta_{5}\leq r_{\max}\,\,.$ (42)
From Equations 28 and 33, we have
$r_{\min}\leq\beta_{1}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}}+\beta_{5}\leq r_{\max}\,\,.$ (43)
Since $\beta_{1}<0$ and $\beta_{5}\leq r_{\max}$, the right side of the above inequality always holds. From the left side, we have
$\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}}\,\,.$ (44)
Since the above equation holds for for all $\frac{O^{A}}{R^{A}}\in(0,c)$, we obtain Item 6:
$\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq c^{\beta_{4}}\,\,.$ (45)
We have now shown all constraints 16 hold. ∎
Appendix E Proof of Theorem 5.3
  • Theorem 5.3 (restated). Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, if $\beta_{1}<0,\beta_{2}+\beta_{4}=0,\beta_{3}=0,0<\beta_{4}\leq 1,r_{\min}<\beta_{5}\leq r_{\max}$ and $\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq c^{\beta_{4}}$, then following items are sufficient for the $c$-Non-Splitting and $c$-smaller-better properties to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
    1. 1.
      $\beta_{1}\beta_{4}(\beta_{4}+1)c^{\beta_{4}-1}(1-c)^{3}\leq-2$
    2. 2.
      $-\beta_{1}\beta_{4}\geq\frac{c^{1-\beta_{4}}}{(1-c)^{2}}$
Proof.
Initially, we expand and simplify the form of the gross amount function according to our assumptions. Then, we prove that Item 1 is sufficient for the $c$-Non-Splitting property to hold. Finally, we prove that Item 2 is sufficient for the $c$-smaller-better property to hold.
Since $\beta_{1}<0,r_{\min}<\beta_{5}$, we have $\beta_{1}(R^{A})^{\beta_{2}}(c^{\textit{AB}}R^{A})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}\leq r_{\max}$. Since $\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq c^{\beta_{4}}$, we have $r_{\min}\leq\beta_{5}+\beta_{1}c^{\beta_{4}}$.
Since $\beta_{2}=-\beta_{4},\beta_{3}=0,\frac{O^{A}}{R^{A}}\leq c$, we have
$\beta_{1}(R^{A})^{\beta_{2}}(c^{\textit{AB}}R^{A})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}=$ $\beta_{1}\times\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}}+\beta_{5}$
$\geq\beta_{5}+\beta_{1}c^{\beta_{4}}$
$\geq r_{\min}$
Then we can expand the trading fee function and the gross amount function as:
$\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\frac{R^{B}}{R^{A}}\times O^{A}\times\left(\beta_{1}\times\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}}+\beta_{5}\right)$
and
$G_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\\ =\frac{R^{B}}{R^{A}}\times O^{A}\times\left(\beta_{1}\times\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}}+\beta_{5}\right)+\frac{R^{B}\times O^{A}}{R^{A}-O^{A}}\,\,.$ (46)
Now we start to prove the $c$-Non-Splitting property holds. We first show that if the $c$-Non-Splitting property holds for $m=2$, then it holds for any $m>2$. Then, we prove that it holds for $m=2$ when Item 1 holds.
Consider the case of $m=2$, where for $O^{A}_{1},O^{A}_{2}>0$, the gross amount of acquiring $O^{A}_{1}+O^{A}_{2}$ is less than the sum of the gross amounts of acquiring $O^{A}_{1}$ and $O^{A}_{2}$:
$G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{1}+O^{A}_{2})<\\ G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{1})+G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{2})\,\,,$ (47)
For $m=3$, we can get that the total gross amount of acquiring $O^{A}_{1}$, $O^{A}_{2}$ and $O^{A}_{3}$ separately is less than the gross amounts of acquiring $O^{A}_{1}+O^{A}_{2}+O^{A}_{3}$ in one time:
$\sum_{j=1}^{3}G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{j})$
$=$ $\left(G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{1})+G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{2})\right)$
$+G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{3})$
[…]
This can be easily generalized to any $m>2$. Therefore, we only need to prove the $c$-Non-Splitting property for $m=2$, which is shown in Equation 47.
Since $G_{SAMM}(R^{A},R^{B},0)=0$, Equation 47 is equivalent to
$G_{SAMM}(R^{A},R^{B},0)+G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{1}+O^{A}_{2})<\\ G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{1})+G_{\textit{SAMM}}(R^{A},R^{B},O^{A}_{2})\,\,.$ (49)
The above inequality holds when $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$ is strictly concave over $O^{A}$. A sufficient condition for strict concavity is the second derivative of $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$ to be negative for all $0<O^{A}<c\times R^{A}$:
$\frac{d^{2}G_{\textit{SAMM}}(R^{A},R^{B},O^{A})}{d\left(O^{A}\right)^{2}}<0$ (50)
From Equation 46, the second derivative of the gross amount function is
$\frac{d^{2}G_{\textit{SAMM}}(R^{A},R^{B},O^{A})}{d\left(O^{A}\right)^{2}}$
$=$ $\beta_{1}\beta_{4}(\beta_{4}+1)\left(R^{A}\right)^{-\beta_{4}-1}R^{B}\left(O^{A}\right)^{\beta_{4}-1}+\frac{2R^{A}R^{B}}{\left(R^{A}-O^{A}\right)^{3}}$
$<$ $\beta_{1}\beta_{4}(\beta_{4}+1)\left(R^{A}\right)^{-\beta_{4}-1}R^{B}\left(O^{A}\right)^{\beta_{4}-1}+\frac{2R^{A}R^{B}}{\left(R^{A}-cR^{A}\right)^{3}}$
$=$ $\beta_{1}\beta_{4}(\beta_{4}+1)\left(R^{A}\right)^{-\beta_{4}-1}R^{B}\left(O^{A}\right)^{\beta_{4}-1}+\frac{2R^{B}}{(1-c)^{3}\left(R^{A}\right)^{2}}\,\,.$
Therefore, a sufficient condition to make the second derivative negative is
$\beta_{1}\beta_{4}(\beta_{4}+1)\left(R^{A}\right)^{-\beta_{4}-1}R^{B}\left(O^{A}\right)^{\beta_{4}-1}+\frac{2R^{B}}{(1-c)^{3}\left(R^{A}\right)^{2}}\leq 0\,\,.$ (51)
The above inequation is equal to
$\beta_{1}\beta_{4}(\beta_{4}+1)(1-c)^{3}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}-1}\leq-2\,\,.$ (52)
The above condition is sufficient for $c$-Non-Splitting property.
Since $\beta_{4}\leq 1$ and $\frac{O^{A}}{R^{A}}<c$, we have
$\beta_{1}\beta_{4}(\beta_{4}+1)(1-c)^{3}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}-1}\leq\\ \beta_{1}\beta_{4}(\beta_{4}+1)c^{\beta_{4}-1}(1-c)^{3}\,\,.$ (53)
Therefore, Item 1 is sufficient for Equation 52, which indicates the $c$-Non-Splitting property holds under Item 1.
Now we turn to the $c$-smaller-better property. The $c$-smaller-better property considers two pools $i$ and $j$, assuming their reported prices are identical. Denote the inverse of this price by $c^{\textit{AB}}$, so $\frac{R^{A}_{i}}{R^{B}_{i}}=\frac{R^{A}_{j}}{R^{B}_{j}}=\frac{1}{c^{\textit{AB}}}>0$. A sufficient condition for the $c$-smaller-better property is the derivative of the gross amount function over $R^{A}$ to be positive for all $0<O^{A}<c\times R^{A}$:
$\frac{dG_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})}{dR^{A}}>0\,\,$ (54)
From Equation 46, we have the derivative:
$\frac{dG_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})}{dR^{A}}=-c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}-\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-O^{A})^{2}}\,\,.$
Since $O^{A}<c\times R^{A}$, we have a lower bound of the derivative:
$\frac{dG_{\textit{SAMM}}(R^{A},c^{\textit{AB}}R^{A},O^{A})}{dR^{A}}>-c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}-\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-cR^{A})^{2}}\,\,.$ (55)
Therefore, it is a sufficient condition for Equation 54 to hold if the lower bound is non-navigate:
$-c^{\textit{AB}}\beta_{1}\beta_{4}\left(\frac{O^{A}}{R^{A}}\right)^{\beta_{4}+1}-\frac{c^{\textit{AB}}(O^{A})^{2}}{(R^{A}-cR^{A})^{2}}\geq 0\,\,.$ (56)
The above equation is equal to
$-\beta_{1}\beta_{4}\geq\frac{1}{(1-c)^{2}}\left(\frac{O^{A}}{R^{A}}\right)^{1-\beta_{4}}\,\,.$ (57)
Since $0<\frac{O^{A}}{R^{A}}<c$ and $\beta_{4}\leq 1$, we have
$\frac{c^{1-\beta_{4}}}{(1-c)^{2}}\geq\frac{1}{(1-c)^{2}}\left(\frac{O^{A}}{R^{A}}\right)^{1-\beta_{4}}\,\,.$ (58)
Therefore, Item 2 is sufficient for Equation 57 to hold, which indicates the $c$-smaller-better property holds under Item 2.
In summary, we have shown that Item 1 is sufficient for the $c$-Non-Splitting property to hold, and Item 2 is sufficient for the $c$-smaller-better property to hold.
Appendix F CPMM Equilibrium
In CPMM, the gross amount is linear to the net amount. Since traders can suffer less from slippage by splitting a trade, the gross amount of splitting a trade is less than the gross amount of trading the same amount at one time. {restatable}thmtheoremuniswapsplitting In $\Gamma_{n}(\textit{tf}_{\textit{CPMM}})$, the following strategy $\tau^{BA}(\textbf{R},b^{\textit{BA}})$ is the only dominant strategy for the BA trader, where
$\tau^{BA}(\textbf{R},b^{\textit{BA}})=\begin{cases}1,&\text{if }a^{\textit{BA}}=\left(\frac{R^{A}_{1}}{\sum R^{A}_{i}}b^{\textit{BA}},\cdots,\frac{R^{A}_{n}}{\sum R^{A}_{i}}b^{\textit{BA}}\right)\\ 0,&\text{Otherwise.}\end{cases}$ (59)
Proof.
We start by calculating the best response for the BA trader. Since the utility $U^{BA}(\textbf{R},b^{\textit{BA}},\pi^{BA})$ is a linear combination of the revenue over actions $U^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ (Equation 15), we can first calculate the optimal action for the BA trader. Combining the CPMM gross amount (Equation 4.3) and the revenue function (Equation 6.1.4), we obtain
$U^{\textit{BA}}(\textbf{R},\ a^{\textit{BA}})$ $=-\sum_{i}\frac{1}{\gamma}\frac{\frac{p^{A}}{p^{B}}R^{A}_{i}b^{\textit{BA}}_{i}}{R^{A}_{i}-b^{\textit{BA}}_{i}}$
$=-\frac{p^{A}}{\gamma p_{B}}\sum_{i}\frac{R^{A}_{i}b^{\textit{BA}}_{i}}{R^{A}_{i}-b^{\textit{BA}}_{i}}\,\,,$ (60)
where the sum of $b^{\textit{BA}}_{i}$ is the total required amount of token A (Equation 11).
We first consider the case of two pools and then extend it to the general case. We define the function $f(\cdot,\cdot)$ which is proportional to the utility of trader in $\textit{pool}_{i}$ and $\textit{pool}_{j}$:
$f(b^{\textit{BA}}_{i},b^{\textit{BA}}_{j})\mathrel{\mathop{:}}=-\left(\frac{R^{A}_{i}b^{\textit{BA}}_{i}}{R^{A}_{i}-b^{\textit{BA}}_{i}}+\frac{R^{A}_{j}b^{\textit{BA}}_{j}}{R^{A}_{j}-b^{\textit{BA}}_{j}}\right)\,\,,$
. Denote by $z\mathrel{\mathop{:}}=b^{\textit{BA}}_{i}+b^{\textit{BA}}_{j}\leq b^{\textit{BA}}$. Then we have
$f(b^{\textit{BA}}_{i},z-b^{\textit{BA}}_{i})=-\left(\frac{R^{A}_{i}b^{\textit{BA}}_{i}}{R^{A}_{i}-b^{\textit{BA}}_{i}}+\frac{R^{A}_{j}(z-b^{\textit{BA}}_{i})}{R^{A}_{j}-(z-b^{\textit{BA}}_{i})}\right)\,\,.$
The derivative of the above function is
$\frac{df(b^{\textit{BA}}_{i},z-b^{\textit{BA}}_{i})}{db^{\textit{BA}}_{i}}=\frac{(R^{A}_{i})^{2}}{(R^{A}_{i}-b^{\textit{BA}}_{i})^{2}}-\frac{(R^{A}_{j})^{2}}{(R^{A}_{j}-(z-b^{\textit{BA}}_{i}))^{2}}\,\,.$
If $0\leq b^{\textit{BA}}_{i}<\frac{R^{A}_{i}}{R^{A}_{i}+R^{A}_{j}}z$, then $\frac{df(b^{\textit{BA}}_{i},z-b^{\textit{BA}}_{i})}{db^{\textit{BA}}_{i}}>0$; if $\frac{R^{A}_{i}}{R^{A}_{i}+R^{A}_{j}}z<b^{\textit{BA}}_{i}\leq z$, then $\frac{df(b^{\textit{BA}}_{i},z-b^{\textit{BA}}_{i})}{db^{\textit{BA}}_{i}}<0$. Therefore, $f(b^{\textit{BA}}_{i},z-b^{\textit{BA}}_{i})$ is maximized only when $b^{\textit{BA}}_{i}=\frac{R^{A}_{i}}{R^{A}_{i}+R^{A}_{j}}z=\frac{R^{A}_{i}}{R^{A}_{i}+R^{A}_{j}}(b^{\textit{BA}}_{i}+b^{\textit{BA}}_{j})$.
The revenue $U^{\textit{BA}}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ reaches the maximum only when $\forall i,j,\frac{b^{\textit{BA}}_{i}}{b^{\textit{BA}}_{j}}=\frac{R^{A}_{i}}{R^{A}_{j}}$, or the trader can replace $b^{\textit{BA}}_{i}$ and $b^{\textit{BA}}_{j}$ with $\frac{R^{A}_{i}}{R^{A}_{i}+R^{A}_{j}}(b^{\textit{BA}}_{i}+b^{\textit{BA}}_{j})$ and $\frac{R^{A}_{j}}{R^{A}_{i}+R^{A}_{j}}(b^{\textit{BA}}_{i}+b^{\textit{BA}}_{j})$ to get higher utility.
Therefore, the only optimal action of the BA trader is
$a^{\textit{BA}}=\left(\frac{R^{A}_{1}}{\sum R^{A}_{i}},\cdots,\frac{R^{A}_{n}}{\sum R^{A}_{i}}\right)\,\,.$ (61)
Then, the only dominant strategy of the BA trader is
$\tau^{BA}(\textbf{R},b^{\textit{BA}})=\begin{cases}1,&\text{if }a^{\textit{BA}}=\left(\frac{R^{A}_{1}}{\sum R^{A}_{i}},\cdots,\frac{R^{A}_{n}}{\sum R^{A}_{i}}\right)\\ 0,&\text{Otherwise.}\end{cases}\qed$ (62)
Appendix G Equilibrium Proof Details
We provide the details for the proofs outlined in Section 7.
G.1 Proof of Lemma 7.1.1
  • Lemma 7.1.1 (restated). In $\Gamma_{n}(tf_{\textit{SAMM}})$, a trader wants to get $b^{\textit{BA}}$ token A when the system state is R. Then for the action which obtains all $b^{\textit{BA}}$ token A in one of the smallest pools with index $i^{\ast}$, where $a_{i^{\ast}}^{\text{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, the trader has no less than the revenue of any other actions:
    $\forall a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}}),U^{\textit{BA}}(a_{i^{\ast}}^{\text{BA}},\textbf{R})\geq U^{\textit{BA}}(a^{\textit{BA}},\textbf{R})\,\,.$
Proof.
Consider an arbitrary action acquiring $b^{\textit{BA}}$ token B, $a^{\textit{BA}}=\left(b^{\textit{BA}}_{1},\cdots,b^{\textit{BA}}_{n}\right)\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})$. Given the state of the pool R=((${\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)}$, the utility of the trader following action $a^{\textit{BA}}$ is
$U^{\textit{BA}}(\textbf{R},a^{\textit{BA}})=-p^{B}\times\sum_{j=1}^{n}G_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p_{B}}R^{A}_{j},b^{\textit{BA}}_{j})\,\,.$ (63)
Since $a_{i^{\ast}}^{\text{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, we have $R^{A}_{i^{\ast}}\leq R^{A}_{j}$. From the $c$-smaller-better property, for all $j$, the gross amount of getting $b^{\textit{BA}}_{j}$ in $pool_{i^{\ast}}$ is no larger than that in $pool_{j}$:
$G_{\textit{SAMM}}(R^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}R^{A}_{i^{\ast}},b^{\textit{BA}}_{j})\leq G_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}}_{j})\,\,.$ (64)
Summing over all $j$, we have
$\sum_{j=1}^{n}G_{\textit{SAMM}}(R^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}R^{A}_{i^{\ast}},b^{\textit{BA}}_{j})\leq\sum_{j=1}^{n}G_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}}_{j})\,\,.$ (65)
From $c$-non-spliting property, since $\sum_{j=1}^{n}b^{\textit{BA}}_{j}=b^{\textit{BA}}$, the gross amount of trading $b^{\textit{BA}}$ in a pool is no larger than the sum of gross amount of trading $b^{\textit{BA}}_{j}$ in the same pool:
$G_{\textit{SAMM}}(R^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}R^{A}_{i^{\ast}},b^{\textit{BA}})\leq\sum_{j=1}^{n}G_{\textit{SAMM}}(R^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}R^{A}_{i^{\ast}},b^{\textit{BA}}_{j})$ (66)
Combining with Equation 65, we obtain
$G_{\textit{SAMM}}(R^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}R^{A}_{i^{\ast}},b^{\textit{BA}})\leq\sum_{j=1}^{n}G_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}}_{j})\,\,.$ (67)
We thus conclude that the revenue of action ${a_{i^{\ast}}^{\text{BA}}(b^{\textit{BA}})=\left(0,\cdots,b^{\textit{BA}}_{i^{\ast}}=b^{\textit{BA}},0,\cdots,0\right)}$ is maximal:
$U^{\textit{BA}}(\textbf{R},a_{i^{\ast}}^{\text{BA}}(b^{\textit{BA}}))=$ $-p^{B}\times G(R^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}R^{A}_{i^{\ast}},b^{\textit{BA}})$
$\geq$ $-p^{B}\times\sum_{j=1}^{n}G_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}}_{j})$
[…]
G.2 Proof of Theorem 7.1.3
  • Theorem 7.1.3 (restated). In $\Gamma_{n}(tf_{\textit{SAMM}})$, considering the following dominant strategy of the BA trader which randomly selects one of the smallest pools to acquire all required tokens:
    $\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}\frac{1}{n_{\min}(\textbf{R})},&\text{if }\exists i,a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\\ 0,&\text{Otherwise.}\end{cases},$ (68)
    then for all strategies $\pi^{BA}$ that have a positive probability of actions not trading in one of the smallest pools, i.e., $\exists a^{\textit{BA}}=\left(b^{\textit{BA}}_{1},\cdots,b^{\textit{BA}}_{i},\cdots,b^{\textit{BA}}_{n}\right)\notin\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R}),\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})>0$ , the utility of the BA trader is strictly lower than with strategy $\tau^{BA}$:
    $U^{\textit{BA}}(\tau^{BA},\textbf{R},b^{\textit{BA}})>U^{\textit{BA}}(\pi^{BA},\textbf{R},b^{\textit{BA}})\,\,.$
Proof.
Denote the minimal amount of deposited token A among all pools by $R^{A}_{\min}=\min_{1\leq i\leq n}R^{A}_{i}$, then all smallest pools have $R^{A}_{\min}$ deposited token A and $\frac{p^{A}}{p^{B}}R^{B}_{\min}$ deposited token B. Then, from Equation 15, the utility of the BA trader under strategy $\tau^{BA}$ is
$U^{\textit{BA}}(\textbf{R},b^{\textit{BA}},\tau^{BA})=\\ \sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})}\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},a^{\textit{BA}})\,\,.$
From the definition of $\tau^{BA}$ (Equation 68), the above equation can be expanded to
$U^{\textit{BA}}(\textbf{R},b^{\textit{BA}},\tau^{BA})=\sum_{a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})}\frac{1}{n_{\min}(\textbf{R})}\times U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})\$ (69)
The revenue of the BA trader under action $a_{i}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})$ is (using Equation 6.1.4)
$U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})=$ $-p^{B}\times\sum_{i=1}^{n}G_{\textit{SAMM}}(R^{A}_{i},R^{B}_{i},b^{\textit{BA}}_{i})$
$=$ $-p^{B}\times G_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{B}_{i},b^{\textit{BA}})\,\,.$
Since $a_{i}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})$, we have $R^{A}_{i}=R^{A}_{\min}$, which means
$U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})=-p^{B}\times G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{B}_{\min},b^{\textit{BA}})\,\,.$ (70)
Combining Equation 69 and 70, we find the utility of the BA trader with strategy $\tau^{BA}$
$U^{\textit{BA}}(\textbf{R},b^{\textit{BA}},\tau^{BA})$
$=$ $\sum_{a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})}\frac{1}{n_{\min}(\textbf{R})}\times\left(-p^{B}\times G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{B}_{\min},b^{\textit{BA}})\right)$
$=$ $-p^{B}\times G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{B}_{\min},b^{\textit{BA}})\,\,.$ (71)
Combining Equation 70, we have
$U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})=U^{\textit{BA}}(\textbf{R},b^{\textit{BA}},\tau^{BA})\,\,.$ (72)
Now, consider the utility of a strategy $\pi^{BA}$. Considering any strategy $\pi^{BA}$ that splits the trade, i.e., $\exists\tilde{a}^{\textit{BA}}=\left(b^{\textit{BA}}_{1},\cdots,b^{\textit{BA}}_{i},\cdots,b^{\textit{BA}}_{n}\right)\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}}),\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})>0,\exists i\neq j,b^{\textit{BA}}_{i}>0,b^{\textit{BA}}_{j}>0$.
Now we consider the revenue of the deviation actions. Considering any strategy $\pi^{BA}$ where $\exists\tilde{a}^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})\setminus\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R}),\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})>0.$ There are two kinds of deviation actions. One is that the trader splits the transaction, namely $\tilde{a}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1}}(b^{\textit{BA}},\textbf{R})$. $\exists i\neq j,b^{\textit{BA}}_{i}>0,b^{\textit{BA}}_{j}>0$. The other is that the trader trades in a non-smallest pool, that is, $\tilde{a}^{\textit{BA}}\in\mathcal{A}^{S}(b^{\textit{BA}})\setminus\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$.
For the first case where $a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})\setminus\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, we have $\exists i\neq j,b^{\textit{BA}}_{i}>0,b^{\textit{BA}}_{j}>0$. The revenue of the BA trader under this action is
$U^{\textit{BA}}(\textbf{R},\tilde{a}^{\textit{BA}})=-p^{B}\times\sum_{i=1}^{n}\left(G_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}}_{i})\right)$ (73)
From the $c$-smaller-better property, since $\forall 1\leq i\leq n,R^{A}_{i}\geq R^{A}_{\min}$, we have
$\sum_{i=1}^{n}\left(G_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}}_{i})\right)\geq\\ \sum_{i=1}^{n}\left(G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{A}_{\min},b^{\textit{BA}}_{i})\right)\,\,.$ (74)
Since $\exists i\neq j,b^{\textit{BA}}_{i}>0,b^{\textit{BA}}_{j}>0$, according to the $c$-non-spliting property, we have
$\sum_{i=1}^{n}\left(G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{A}_{\min},b^{\textit{BA}}_{i})\right)>G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{A}_{\min},b^{\textit{BA}})\,\,.$
Therefore, $\forall\tilde{a}^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})\setminus\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, the revenue of the action $\tilde{a}^{\textit{BA}}$ (Equation 73) can be expanded as
$U^{\textit{BA}}(\textbf{R},\tilde{a}^{\textit{BA}})$ $=-p^{B}\times\sum_{i=1}^{n}\left(G_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}}_{i})\right)$
$<$ $-p^{B}\times G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{A}_{\min},b^{\textit{BA}})\,\,.$ (75)
Since the right part of the above inequality is the revenue of the action $a_{i}^{\textit{BA}}$ (Equation 70), the revenue of action $\tilde{a}^{\textit{BA}}$ is strictly smaller than the revenue of action $a_{i}^{\textit{BA}}$:
$U^{\textit{BA}}(\textbf{R},\tilde{a}^{\textit{BA}})<U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})\,\,.$ (76)
Now we turn to the second case of an action that acquiring all tokens in a non-smallest pool, i.e., $\tilde{a}^{\textit{BA}}\in\mathcal{A}^{S}(b^{\textit{BA}})\setminus\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$. Here, we rewrite $\tilde{a}^{\textit{BA}}$ as $\tilde{a}_{j}^{\textit{BA}}$, the action acquiring all $b^{\textit{BA}}$ in $\textit{pool}_{j}$, where $R^{A}_{j}>R^{A}_{\min}$. Then, the revenue of the BA trader is
$U^{\textit{BA}}(\textbf{R},\tilde{a}_{j}^{\textit{BA}})=-p^{B}\times G_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}})\,\,.$ (77)
From the $c$-smaller-better property, since $R^{A}_{j}>R^{A}_{\min}$, we have
$G_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}})<G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{A}_{\min},b^{\textit{BA}})\,\,.$
Therefore, from Equation 77, we have
$U^{\textit{BA}}(\textbf{R},\tilde{a}_{j}^{\textit{BA}})<-p^{B}\times G_{\textit{SAMM}}(R^{A}_{\min},\frac{p^{A}}{p^{B}}R^{A}_{\min},b^{\textit{BA}})$
Since the right part of the above inequality is the utility of the action $a_{i}^{\textit{BA}}$ (Equation 70), the revenue of the action $\tilde{a}_{j}^{\textit{BA}}$ is strictly lower than the revenue of the action $a_{i}^{\textit{BA}}$ when $\tilde{a}_{j}^{\textit{BA}}\in\mathcal{A}^{S}(b^{\textit{BA}})\setminus\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$:
$U^{\textit{BA}}(\textbf{R},\tilde{a}_{j}^{\textit{BA}})<U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})\,\,.$ (78)
Combining the conditions for Equation 76 and 78, then $\forall\tilde{a}^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})\setminus\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, we have
$U^{\textit{BA}}(\textbf{R},\tilde{a}^{\textit{BA}})<U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})\,\,.$ (79)
Now we return to the utility of the BA trader with strategy $\pi^{BA}$. We tease out the deviating action:
$U^{\textit{BA}}(\textbf{R},b^{\textit{BA}},\pi^{BA})$
$=$ $\sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})}\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},a^{\textit{BA}})$
$=$ $\sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})\setminus\{\tilde{a}^{\textit{BA}}\}}\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},a^{\textit{BA}})$
$+$ $\pi^{BA}(\textbf{R},b^{\textit{BA}},\tilde{a}^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},\tilde{a}^{\textit{BA}})\,\,.$ (80)
From lemma 7.1.1, the revenue of any action ${a}_{t}^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})$ is no larger than the revenue of the action $a_{i}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}})$. Combining Equation 70, we have
$\sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})\setminus\{\tilde{a}^{\textit{BA}}\}}\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},a^{\textit{BA}})\leq\\ \sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})\setminus\{\tilde{a}^{\textit{BA}}\}}\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})$ (81)
Combining Equations 79 anc 81, we expand the utility of the BA trader under strategy $\pi^{BA}$ in Equation G.2 as
$U^{\textit{BA}}(\textbf{R},b^{\textit{BA}},\pi^{BA})$
$=$ $\sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})-\{\tilde{a}^{\textit{BA}}\}}\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},a^{\textit{BA}})$
$+$ $\pi^{BA}(\textbf{R},b^{\textit{BA}},\tilde{a}^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},\tilde{a}^{\textit{BA}})$
[…]
$+$ $\pi^{BA}(\textbf{R},b^{\textit{BA}},\tilde{a}^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},\tilde{a}^{\textit{BA}})$
[…]
$+$ $\pi^{BA}(\textbf{R},b^{\textit{BA}},\tilde{a}^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})$
$=$ $\sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})}\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})\times U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})$
$=$ $U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})\times\sum_{a^{\textit{BA}}\in\mathcal{A}^{\textit{BA}}(b^{\textit{BA}})}\pi^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$
$=$ $U^{\textit{BA}}(\textbf{R},a_{i}^{\textit{BA}})$
[…]
The above inequality indicates that the utility of the BA trader under strategy $\pi^{BA}$ is strictly lower than the utility of the BA trader under strategy $\tau^{BA}$. ∎
G.3 Proof of Lemma 7.2.1
  • Lemma 7.2.1 (restated). For any action of liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$, if $\rho^{A}(\textbf{R}+a_{lp})\geq\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})$, then $a_{lp}=a_{lp}^{\textit{fill}}$.
Proof.
Consider any $j$ s.t. $l^{A^{\ast}}_{j}>0$, from the definition of $a_{lp}^{\textit{fill}}$ (Definition 7.2.1), $l^{A^{\ast}}_{j}+R^{A}_{j}$ is the minimal among all pools with state $\textbf{R}+a_{lp}$, i.e.,
$\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})=\hat{l}^{A}_{j}+R^{A}_{j}\,\,.$ (82)
Since then all pools in $\textbf{R}+a_{lp}$ have no less than $\rho^{A}(\textbf{R}+a_{lp})$ deposited token A, if $\rho^{A}(\textbf{R}+a_{lp})\geq\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})$, we have
$l^{A}_{j}+R^{A}_{j}\geq\rho^{A}(\textbf{R}+a_{lp})\geq\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})=\hat{l}^{A}_{j}+R^{A}_{j}\,\,.$ (83)
Therefore, we have
$l^{A}_{j}\geq\hat{l}^{A}_{j}\,\,.$ (84)
Considering the sum of $l^{A}_{j}$ and $\hat{l}^{A}_{j}$, we have
$\sum_{l^{A}_{j}>0}l^{A}_{i}\geq\sum_{\hat{l}^{A}_{j}>0}\hat{l}^{A}_{j}=l^{A}\,\,.$ (85)
For $a_{lp}$, the sum of input tokens in all pools is $l^{A}$:
$\sum_{i=1}^{n}l^{A}_{i}=l^{A}\,\,.$ (86)
And $\forall 1\leq i\leq n$, $l^{A}_{i}$ is non-negative,
$l^{A}_{i}\geq 0\,\,.$ (87)
Combining Expressions 84 8586 and 87, all inequation holds with equality, which indicates that $a_{lp}=a_{lp}^{\textit{fill}}$:
$\forall 1\leq i\leq n,l^{A}_{i}=\hat{l}^{A}_{i}\,\,.\qed$ (88)
G.4 Proof of Lemma 7.2.1
  • Lemma 7.2.1 (restated). For any two pools $i$ and $j$, if $R^{A}_{i}<R^{A}_{j}$, for any output amount $b^{\textit{BA}}$ of token B, the trading fee of $\textit{pool}_{i}$ is strictly smaller than the trading fee of $\textit{pool}_{j}$:
    $tf_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}})<tf_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}})\,\,.$
Proof.
From $c$-smaller-better property, since $R^{A}_{i}<R^{A}_{j}$, the gross amount of $\textit{pool}_{i}$ is strictly smaller than the gross amount of $\textit{pool}_{j}$,
$G_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}})<G_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}})\,\,.$
Since the gross amount is the sum of the net amount and the trading fee, by expanding the above inequality, we have
$tf_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}})+\textit{net}^{B}(R^{A},\frac{p^{A}}{p^{B}}R^{A}_{i},O^{A})<\\ tf_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}})+\textit{net}^{B}(R^{A},\frac{p^{A}}{p^{B}}R^{A}_{j},O^{A})\,\,.$ (89)
Considering the net amount of these two pools, since $R^{A}_{i}<R^{A}_{j}$, $\textit{pool}_{i}$ has higher slippage than $\textit{pool}_{j}$:
$\textit{net}^{B}(R^{A},\frac{p^{A}}{p^{B}}R^{A}_{i},O^{A})>\textit{net}^{B}(R^{A},\frac{p^{A}}{p^{B}}R^{A}_{j},O^{A})\,\,.$ (90)
Combining Equation 89 and 90, we conclude that the trading fee of $\textit{pool}_{i}$ is strictly smaller than the trading fee of $\textit{pool}_{j}$:
$tf_{\textit{SAMM}}(R^{A}_{i},\frac{p^{A}}{p^{B}}R^{A}_{i},b^{\textit{BA}})<tf_{\textit{SAMM}}(R^{A}_{j},\frac{p^{A}}{p^{B}}R^{A}_{j},b^{\textit{BA}})\,\,.\qed$ (91)
G.5 Proof of Theorem 7.2.2
  • Theorem 7.2.2 (restated). Denote by $\hat{a}_{lp}=\left(\left(\frac{1}{n}l^{A},\frac{1}{n}l^{B}\right),\cdots\right)$ the action of evenly depositing tokens in all pools. In $\Gamma_{n}(tf_{\textit{SAMM}})$, if for all $i$ and $j$ that the liquidity amounts are the same, $R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, the liquidity provider strategy which only takes action $\hat{a}_{lp}$,
    $\tau_{lp}(\textbf{R},l^{A},l^{B},a_{lp})=\begin{cases}1,&\text{if }a_{lp}=\hat{a}_{lp}\\ 0,&\text{Otherwise.}\end{cases}\$
    and any best response of the trader constitutes an SPNE.
Proof.
Without loss of generality, we only consider BA traders since actions of AB traders are symmetric.
Since the traders have given the best response, we only need to prove that $\tau_{lp}$ is also the best response.
Consider any action of liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$, the pool state after this action is $\textbf{R}+a_{lp}$. From Corollary 7.1.3, any best response of the trader only has a positive probability of taking an action that trades in exactly one of the smallest pools. Therefore, from Equation 18, when the trader use any best response $\tau^{BA}$, the liquidity provider’s revenue with action $a_{lp}$ is
$U_{lp}(\textbf{R},l^{A},l^{B},a_{lp},\tau^{BA},\tau^{AB})=\\ E_{b^{\textit{BA}}\sim D^{BA}}\left[\sum_{\begin{subarray}{c}a_{i}^{\textit{BA}}(b^{\textit{BA}})\\ \in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R}+a_{lp})\end{subarray}}\left(\begin{subarray}{c}\tau^{BA}(\textbf{R}+a_{lp},b^{\textit{BA}},a_{i}^{\textit{BA}}(b^{\textit{BA}}))\times\\ U_{lp}(\textbf{R},a_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))\end{subarray}\right)\right]$ (92)
Since $\forall 1\leq i,j\leq n$, $R^{A}_{i}=R^{A}_{j}$. Therefore, the minimal deposited amount of token A among all pools in $\textbf{R}+a_{lp}$ is:
$\rho^{A}(\textbf{R}+a_{lp})=R^{A}_{1}+l^{A}_{\min}\,\,.$ (93)
From Equation 6.1.4, the revenue of a liquidity provider due to action $a_{lp}$ and the BA trader action $a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R}+a_{lp})$ is acquiring all $b^{\textit{BA}}$ in one of the smallest pool after the liquidity provider’s action, say $\textit{pool}_{i}$, is
$U_{lp}(\textbf{R},a_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))$
$=$ $p^{B}\times\textit{tf}(R^{A}_{i}+l^{A}_{i},\frac{p^{A}}{p^{B}}\left(R^{A}_{i}+l^{A}_{i}\right),b^{\textit{BA}})\times\frac{l^{A}_{i}}{l^{A}_{i}+R^{A}_{i}}$
$=$ $p^{B}\times\textit{tf}(R^{A}_{1}+l^{A}_{\min},\frac{p^{A}}{p^{B}}\left(R^{A}_{1}+l^{A}_{\min}\right),b^{\textit{BA}})\times\frac{l^{A}_{\min}}{l^{A}_{\min}+R^{A}_{1}}\,\,.$ (94)
When $a_{lp}=\hat{a}_{lp}$, the liquidity provider inputs tokens in all pools evenly, i.e., $l^{A}_{i}=\frac{1}{n}l^{A}$. Then each pool is identical, the trader’s choice of $\textit{pool}_{i}$ has the same revenue for the liquidity provider as $\textit{pool}_{1}$:
$U_{lp}(\textbf{R},\hat{a}_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))=U_{lp}(\textbf{R},\hat{a}_{lp},a_{1}^{\textit{BA}}(b^{\textit{BA}}))$
Therefore, from Equation 92, the revenue of the liquidity provider under action $\hat{a}_{lp}$ is
$U_{lp}(\textbf{R},l^{A},l^{B},\hat{a}_{lp},\tau^{BA},\tau^{AB})=\\ E_{b^{\textit{BA}}\sim D^{BA}}\left[U_{lp}(\textbf{R},\hat{a}_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))\right]$ (95)
Since $\sum_{i=1}^{n}l^{A}_{i}=l^{A}$, when $\forall 1\leq i\leq n,l^{A}_{i}=\frac{1}{n}l^{A}$, we have for any action $a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B}),l^{A}_{\min}\leq\frac{1}{n}l^{A}$. Then,
$\frac{l^{A}_{\min}}{l^{A}_{\min}+R^{A}_{1}}\leq\frac{\frac{1}{n}l^{A}}{\frac{1}{n}l^{A}+R^{A}_{1}}\,\,.$ (96)
From Lemma 7.2.1, the trading fee of the smallest pool is no larger than the trading fee of any other pool:
$tf_{\textit{SAMM}}(R^{A}_{1}+l^{A}_{\min},\frac{p^{A}}{p^{B}}\left(R^{A}_{1}+l^{A}_{\min}\right),b^{\textit{BA}})\leq\\ tf_{\textit{SAMM}}(R^{A}_{i}+\frac{1}{n}l^{A},\frac{p^{A}}{p^{B}}\left(R^{A}_{i}+\frac{1}{n}l^{A}\right),b^{\textit{BA}})\,\,.$ (97)
Combining Equation 96 and 97, we have
$p^{B}\times\textit{tf}(R^{A}_{1}+l^{A}_{\min},\frac{p^{A}}{p^{B}}\left(R^{A}_{1}+l^{A}_{\min}\right),b^{\textit{BA}})\times\frac{l^{A}_{\min}}{l^{A}_{\min}+R^{A}_{1}}\leq\\ p^{B}\times\textit{tf}(R^{A}_{i}+\frac{1}{n}l^{A},\frac{p^{A}}{p^{B}}\left(R^{A}_{i}+\frac{1}{n}l^{A}\right),b^{\textit{BA}})\times\frac{\frac{1}{n}l^{A}}{\frac{1}{n}l^{A}+R^{A}_{1}}\,\,,$ (98)
which indicates that the revenue of the liquidity provider under action $\hat{a}_{lp}$ is not smaller than any other action when traders take action $a_{i}^{\textit{BA}}(b^{\textit{BA}})$, i.e.,
$U_{lp}(\textbf{R},a_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))\leq U_{lp}(\textbf{R},\hat{a}_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))\,\,.$ (99)
Combining Equation 92 and 99, the revenue of the liquidity provider under action $\hat{a}_{lp}$ is not smaller than the revenue of the liquidity provider under any other action when traders use any best response $\tau^{BA}$.
$U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})$
[…]
Consider the utility of the liquidity provider under action $\hat{a}_{lp}$ of evenly depositing tokens in all pools, system state $\textbf{R},l^{A},l^{B}$, and the BA trader strategy $\tau^{BA}$. From the definition of the utility function of the liquidity provider (Equation 19), the utility of the liquidity provider’s strategy $\tau_{lp}$ is equal to the revenue of the liquidity provider under action $\hat{a}_{lp}$:
$U_{lp}(\textbf{R},l^{A},l^{B},\tau_{lp},\tau^{BA},\tau^{AB})$
$=$ $\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\left(\tau_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\times U_{lp}(\textbf{R},l^{A},l^{B},a_{lp},\tau^{BA},\tau^{AB})\right)$
$=$ $U_{lp}(\textbf{R},l^{A},l^{B},\hat{a}_{lp},\tau^{BA},\tau^{AB})$ (101)
Then the liquidity provider strategy $\tau_{lp}$ has no smaller utility than $\pi^{BA}$:
$U_{lp}(\textbf{R},l^{A},l^{B},\pi_{lp},\tau^{BA},\tau^{AB})$
[…]
$=$ $U_{lp}(\textbf{R},l^{A},l^{B},\hat{a}_{lp},\tau^{BA},\tau^{AB})$
[…]
That is, the liquidity provider strategy $\tau_{lp}$ is the best response to the trader strategy $\tau^{BA}$.
Therefore, the liquidity provider strategy $\tau_{lp}$ and any best response of the BA trader $\tau^{BA}$ are an SPNE. ∎
G.6 Proof of Theorem 7.2.3
  • Theorem 7.2.3 (restated). In $\Gamma_{n}(\textit{tf}_{\textit{SAMM}})$, in all SPNE, the liquidity provider’s best response is the fillup strategy:
    $\tau^{\textit{fill}}_{lp}(\textbf{R},l^{A},l^{B},a_{lp})=\begin{cases}1,&\text{if }a_{lp}=a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})\\ 0,&\text{Otherwise.}\end{cases}\,\,.$
Proof.
We prove this by contradiction. (1)Initially, for any action that is not the fillup action, we construct another action resulting in a larger smallest pool. (2)Second, we prove that the constructed action has higher utility than the original action. (3)Third, for any strategy that does not always take the fillup action, we construct a new strategy based on the constructed action. (4)Finally, we prove that the new strategy has a higher utility than the original strategy, which means that the original strategy is not the best response.
(1) Construction of the new action: Consider any action of a liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$. We construct another action $a_{lp}^{\prime}$. First, we want the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ to be larger than the smallest pool in $\textbf{R}+a_{lp}$, which lead to a higher trading fee in a single trade according to Lemma 7.2.1. Second, we want the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ to be unique and also the smallest in R to make the liquidity provider have more shares in the smallest pool than in $\textbf{R}+a_{lp}$.
Denote by $i^{\ast}$ the smallest pool in R with the smaller index, i.e.
$\forall j,R^{A}_{i^{\ast}}\leq R^{A}_{j}$
$\forall j,R^{A}_{i^{\ast}}=R^{A}_{j}\Rightarrow i^{\ast}\leq j\,\,.$ (102)
Consider the fillup action $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})=\left(\left(\hat{l}^{A}_{1},\hat{l}^{B}_{1}\right),\cdots,\left(\hat{l}^{A}_{n},\hat{l}^{B}_{n}\right)\right)$. If $a_{lp}=a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$, we take $a_{lp}^{\prime}=a_{lp}$ and we are done. If $a_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$, from Lemma 7.2.1, we have $\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp})<\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}))$. Denote this difference by $z=\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}))-\rho^{A}(\textbf{R},l^{A},l^{B}+a_{lp})$. We construct the action $a_{lp}^{\prime}=\left(\left(l^{A^{\prime}}_{1},l^{B^{\prime}}_{1}\right),\cdots,\left(l^{A^{\prime}}_{n},l^{B^{\prime}}_{n}\right)\right)$ as follows. Figure 2 shows an example of the construction.
[…]
Then, $\textit{pool}_{i^{\ast}}$ in $\textbf{R}+a_{lp}^{\prime}$ is the only pool with the minimal amount of deposited token A, make it the best choice for the subsequent trader.
$\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R}+a_{lp}^{\prime})=\{a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}})\}\,\,.$ (104)
From the Definition 7.2.1, $\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})=l^{A^{\ast}}_{i^{\ast}}+R^{A}_{i^{\ast}}$. Then, the smallest amount of deposited token A in $\textbf{R}+a_{lp}^{\prime}$ is larger than that in $\textbf{R}+a_{lp}$:
$l^{A^{\prime}}_{i^{\ast}}+R^{A}_{i^{\ast}}$ $=\hat{l}^{A}_{i^{\ast}}-\frac{1}{2}z+R^{A}_{i^{\ast}}$
$=(\hat{l}^{A}_{i^{\ast}}+R^{A}_{i^{\ast}})-\frac{1}{2}z$
$=\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})-\frac{1}{2}z$
$>\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})-z$
$=\rho^{A}(\textbf{R}+a_{lp})\,\,.$ (105)
(2) The revenues of actions $a_{lp}$ and $a_{lp}^{\prime}$: To compare the revenue due to both actions, we consider the strategies and utility of the subsequent trader. Any best response of the trader $\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ uses the smallest pool (Lemma 7.1.3):
$\sum_{a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})}\tau^{BA}(\textbf{R},b^{\textit{BA}},a_{i}^{\textit{BA}}(b^{\textit{BA}}))=1\,\,.$ (106)
For the pool state $\textbf{R}+a_{lp}^{\prime}$, combining Equation 104, the probability of taking action $a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}})$ is 1:
$\tau^{BA}(\textbf{R},b^{\textit{BA}},a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}}))$
$=$ $\sum_{a_{i}^{\textit{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})}\tau^{BA}(\textbf{R},b^{\textit{BA}},a_{i}^{\textit{BA}}(b^{\textit{BA}}))$
$=$ $1\,\,.$ (107)
Therefore, from Equation 18, the utility of the liquidity provider following action $a_{lp}^{\prime}$ and trader’s best response strategy $\tau^{BA}$ is
$U_{lp}(\textbf{R},a_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})=\\ E_{b^{\textit{BA}}\sim D^{BA}}\left[U_{lp}(\textbf{R},a_{lp}^{\prime},a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}}))\right].$ (108)
Similarly, the utility following action $a_{lp}$ is
$U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})=\\ E_{b^{\textit{BA}}\sim D^{BA}}\left[\sum_{\begin{subarray}{c}a_{i}^{\textit{BA}}(b^{\textit{BA}})\\ \in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\end{subarray}}\left(\begin{subarray}{c}U_{lp}(\textbf{R},a_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))\\ \times\tau^{BA}(\textbf{R}+a_{lp},b^{\textit{BA}},a_{i}^{\textit{BA}})\end{subarray}\right)\right].$ (109)
Thus, the revenue following action $a_{lp}^{\prime}$ is (using Equation 6.1.4):
$U_{lp}(\textbf{R},a_{lp}^{\prime},a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}}))=\\ p^{B}\times\textit{tf}(R^{A}_{i^{\ast}}+l^{A^{\prime}}_{i^{\ast}},\frac{p^{A}}{p^{B}}(R^{A}_{i^{\ast}}+l^{A^{\prime}}_{i^{\ast}}),b^{\textit{BA}})\times\frac{l^{A^{\prime}}_{i^{\ast}}}{l^{A^{\prime}}_{i^{\ast}}+R^{A}_{i^{\ast}}}\,\,.$ (110)
Similarly, we can expand the expression of $U_{lp}(\textbf{R},a_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))$ for $a_{i}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R}+a_{lp})$ as
$U_{lp}(\textbf{R},a_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))=\\ p^{B}\times\textit{tf}(R^{A}_{i}+l^{A}_{i},\frac{p^{A}}{p^{B}}(R^{A}_{i}+l^{A}_{i}),b^{\textit{BA}})\times\frac{l^{A}_{i}}{l^{A}_{i}+R^{A}_{i}}\,\,.$ (111)
Since $a_{i}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R}+a_{lp})$, $\textit{pool}_{i}$ has the smallest amount of deposited token A in $\textbf{R}+a_{lp}$:
$R^{A}_{i}+l^{A}_{i}=\rho^{A}(\textbf{R}+a_{lp})\,\,.$ (112)
Interpreting Equation G.6, the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ is larger than the smallest pool in $\textbf{R}+a_{lp}$:
$l^{A^{\prime}}_{i^{\ast}}+R^{A}_{i^{\ast}}>R^{A}_{i}+l^{A}_{i}\,\,.$ (113)
From lemma 7.2.1, the trading fee of trading in a larger pool is larger,
$\textit{tf}(R^{A}_{i^{\ast}}+l^{A^{\prime}}_{i^{\ast}},\frac{p^{A}}{p^{B}}(R^{A}_{i^{\ast}}+l^{A^{\prime}}_{i^{\ast}}),b^{\textit{BA}})>\textit{tf}(R^{A}_{i}+l^{A}_{i},\frac{p^{A}}{p^{B}}(R^{A}_{i}+l^{A}_{i}),b^{\textit{BA}})\,\,.$ (114)
From the definition of $i^{\ast}$ in Equation G.6, we have
$R^{A}_{i^{\ast}}\leq R^{A}_{i}\,\,.$ (115)
Combine Equations 113 and 115, we have
$\frac{l^{A^{\prime}}_{i^{\ast}}}{l^{A^{\prime}}_{i^{\ast}}+R^{A}_{i^{\ast}}}=1-\frac{R^{A}_{i^{\ast}}}{l^{A^{\prime}}_{i^{\ast}}+R^{A}_{i^{\ast}}}>1-\frac{R^{A}_{i}}{l^{A}_{i}+R^{A}_{i}}=\frac{l^{A}_{i}}{l^{A}_{i}+R^{A}_{i}}\,\,.$ (116)
Combining Equations 110111114 and 116, the revenue of the liquidity provider with action $a_{lp}^{\prime}$ is higher than with action $a_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$:
$U_{lp}(\textbf{R},a_{lp}^{\prime},a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}}))>U_{lp}(\textbf{R},a_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))\,\,.$ (117)
Combining Equations 108 and 109, the utility of the liquidity provider under action $a_{lp}^{\prime}$ is higher than that under action $a_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$:
$U_{lp}(\textbf{R},a_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})>U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$ (118)
When $a_{lp}=a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$, we have $a_{lp}^{\prime}=a_{lp}$. Therefore, the following inequality holds for all $a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})$:
$U_{lp}(\textbf{R},a_{lp}^{\prime},a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}}))\geq U_{lp}(\textbf{R},a_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))\,\,.$ (119)
(3) Construction of a new strategy: We showed that for any action $a_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$, the constructed action $a_{lp}^{\prime}$ has a higher revenue. Now, for any liquidity provider strategy $\pi_{lp}$, we can construct a new strategy $\pi_{lp}^{\prime}$ where for every action $a_{lp}$, the probability of constructed action $a_{lp}^{\prime}$ in the new strategy is the same as the original action $a_{lp}$ in the original strategy:
$\pi_{lp}^{\prime}(\textbf{R},l^{A},l^{B},a_{lp}^{\prime})=\pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\,\,.$
If the original strategy is different from the fillup strategy $\pi_{lp}\neq\tau^{\textit{fill}}_{lp}$, then $\exists\tilde{a}_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})$, s.t. $\tilde{a}_{lp}\neq a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ and $\pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})>0$.
(4) Comparison of utilities: From the definition (Equation 19), the utility of the liquidity provider under $\pi_{lp}$ is
$U_{lp}(\textbf{R},\pi_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})=\\ \sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\times U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$ (120)
Similarly, the utility of the liquidity provider under $\pi_{lp}^{\prime}$ is
$U_{lp}(\textbf{R},\pi_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})$
$=$ $\sum_{a_{lp}^{\prime}\in\mathcal{A}_{lp}(l^{A},l^{B})}\pi_{lp}^{\prime}(\textbf{R},l^{A},l^{B},a_{lp}^{\prime})\times U_{lp}(\textbf{R},a_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})$
$=$ $\sum_{a_{lp}^{\prime}\in\mathcal{A}_{lp}(l^{A},l^{B})\setminus\{\tilde{a}_{lp}^{\prime}\}}\left(\begin{subarray}{c}\pi_{lp}^{\prime}(\textbf{R},l^{A},l^{B},a_{lp}^{\prime})\\ \times U_{lp}(\textbf{R},a_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})\end{subarray}\right)$
$+\pi_{lp}^{\prime}(\textbf{R},l^{A},l^{B},\tilde{a}_{lp}^{\prime})\times U_{lp}(\textbf{R},\tilde{a}_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})$ (121)
Since $\pi_{lp}^{\prime}(\textbf{R},l^{A},l^{B},\tilde{a}_{lp}^{\prime})=\pi_{lp}(\textbf{R},l^{A},l^{B},\tilde{a}_{lp})>0$, from Equation 117, we have
$\pi_{lp}^{\prime}(\textbf{R},l^{A},l^{B},\tilde{a}_{lp}^{\prime})\times U_{lp}(\textbf{R},\tilde{a}_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})>\\ \pi_{lp}(\textbf{R},l^{A},l^{B},\tilde{a}_{lp})\times U_{lp}(\textbf{R},\tilde{a}_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$ (122)
Since $\pi_{lp}^{\prime}(\textbf{R},l^{A},l^{B},a_{lp}^{\prime})=\pi_{lp}(\textbf{R},l^{A},l^{B},{a}_{lp})>0$, from Equation 119, we have
$\pi_{lp}^{\prime}(\textbf{R},l^{A},l^{B},a_{lp}^{\prime})\times U_{lp}(\textbf{R},a_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})\geq\\ \pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\times U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$ (123)
Combining Equations 120  G.6122 and 123, the utility of the liquidity provider under $\pi_{lp}^{\prime}$ is higher than that under $\pi_{lp}$:
$U_{lp}(\textbf{R},\pi_{lp}^{\prime},\tau^{BA},\tau^{AB},l^{A},l^{B})>U_{lp}(\textbf{R},\pi_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$
Therefore, any liquidity provider strategy $\pi_{lp}\neq\tau^{\textit{fill}}_{lp}$ is not the best response when the trader follows any best response. Therefore, in all SPNE, the liquidity provider’s best response is the fillup strategy $\tau^{\textit{fill}}_{lp}$. ∎
G.7 Proof of Theorem 7.2.4
  • Theorem 7.2.4 (restated). In $\Gamma_{n}(\textit{tf}_{SAMM})$, assume that the pool state in step $k$ is $\textbf{R}({k})$, $i^{\ast}=i_{\min}(\textbf{R}({k}))$ is the index of the pool with the smallest amount of deposited token A in $\textbf{R}({k})$. Then the trader strategy is to trade in the smallest pool in the last step if it is the smallest one, or randomly select one of the smallest pools, namely,
    $\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}1,&\text{if}R_{i^{\ast}}=\rho^{A}(\textbf{R}({k}),l^{A},l^{B})\text{ and }\\ &a^{\textit{BA}}=a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}})\\ \frac{1}{n_{\min}(\textbf{R})},&\text{if }R_{i^{\ast}}=\rho^{A}(\textbf{R}({k}),l^{A},l^{B})\text{ and }\\ &a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\\ 0,&\text{Otherwise.}\end{cases}$ (124)
    and the liquidity provider’s fillup strategy:
    $\tau^{\textit{fill}}_{lp}(\textbf{R},l^{A},l^{B},a_{lp})=\begin{cases}1,&\text{if }a_{lp}=a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})\\ 0,&\text{Otherwise.}\end{cases},$
    are an SPNE in step $k$.
Proof.
Considering the revenue of the liquidity provider under the action $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$. The amount of deposited token A of $\textit{pool}_{i^{\ast}}$ in $\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ is $R^{A}_{i}+\hat{l}^{A}_{i^{\ast}}$.
We first prove that $\hat{l}^{A}_{i^{\ast}}>0$ by contradiction. If $\hat{l}^{A}_{i^{\ast}}=0$, then $R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}}=R^{A}_{i^{\ast}}$. Then for any input amount $\hat{l}^{A}_{j}>0$, we have $R^{A}_{j}+\hat{l}^{A}_{j}>R^{A}_{j}$. From the definition of $i^{\ast}=i_{\min}(\textbf{R})$ (Equation 7.2.4), we have $R^{A}_{i^{\ast}}\leq R^{A}_{j}$. Therefore, the amount of deposited token A of $\textit{pool}_{i^{\ast}}$ in $\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ is strictly smaller than $\textit{pool}_{j}$
$R^{A}_{j}+\hat{l}^{A}_{j}>R^{A}_{j}\geq R^{A}_{i^{\ast}}=R^{A}_{j}+\hat{l}^{A}_{i^{\ast}}\,\,.$ (125)
This contradicts the definition of $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ (Definition 7.2.1) since $\hat{l}^{A}_{j}>0$. Therefore, $\hat{l}^{A}_{i^{\ast}}>0$. Then also from that definition, the amount of deposited token A of $\textit{pool}_{i^{\ast}}$ after the fillup action is smallest among all pools:
$R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}}=\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}))\,\,.$ (126)
Therefore, from the definition of $\tau^{BA}$ (Equation 124), we have $\tau^{BA}(\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}),b^{\textit{BA}},a_{i^{\ast}}^{\textit{BA}})=1$.
Then, we turn to the revenue of the liquidity provider with the fillup action and prove that it is no smaller than any other action. From Equation 18, the utility of the liquidity provider under the fill-up action and the trader’s best response $\tau^{BA}$ is:
$U_{lp}(\textbf{R},a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}),\tau^{BA},\tau^{AB},l^{A},l^{B})=\\ E_{b^{\textit{BA}}\sim D^{BA}}\left[U_{lp}(\textbf{R},a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}),a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}}))\right]\,\,.$ (127)
Similarly, the utility of the liquidity provider under any action $a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})$ and the trader’s best response $\tau^{BA}$ is
$U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})=\\ E_{b^{\textit{BA}}\sim D^{BA}}\left[\sum_{\begin{subarray}{c}a_{i}^{\textit{BA}}(b^{\textit{BA}})\\ \in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\end{subarray}}\left(\begin{subarray}{c}U_{lp}(\textbf{R},a_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))\\ \times\tau^{BA}(\textbf{R}+a_{lp},b^{\textit{BA}},a_{i}^{\textit{BA}})\end{subarray}\right)\right]\,\,.$ (128)
From the definition of $U_{lp}(\textbf{R},a_{lp},a^{\textit{BA}})$ (Equation 6.1.4), we have
$U_{lp}(\textbf{R},a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}),a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}}))=\\ p^{B}\times\textit{tf}(R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}(R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}}),b^{\textit{BA}})\times\frac{\hat{l}^{A}_{i^{\ast}}}{\hat{l}^{A}_{i^{\ast}}+R^{A}_{i^{\ast}}}\,\,,$ (129)
and
$U_{lp}(\textbf{R},a_{lp},a_{i}^{\textit{BA}}(b^{\textit{BA}}))=\\ p^{B}\times\textit{tf}(R^{A}_{i}+l^{A}_{i},\frac{p^{A}}{p^{B}}(R^{A}_{i}+l^{A}_{i}),b^{\textit{BA}})\times\frac{l^{A}_{i}}{l^{A}_{i}+R^{A}_{i}}\,\,.$ (130)
Since $a_{i}^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, $\textit{pool}_{i}$ has the smallest amount of deposited token A in $\textbf{R}+a_{lp}$:
$R^{A}_{i}+l^{A}_{i}=\rho^{A}(\textbf{R}+a_{lp})\,\,.$ (131)
From Lemma 7.2.1, the minimal amount of deposited token A in $\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ is no less than that in $\textbf{R}+a_{lp}$ for any $a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})$:
$\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}))\geq\rho^{A}(\textbf{R}+a_{lp})\,\,.$ (132)
Combining Equations 126131 and 132, we have
$R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}}\geq R^{A}_{i}+l^{A}_{i}\,\,.$ (133)
From Lemma 7.2.1, the trading fee of trading in a larger pool is larger:
$\textit{tf}(R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}(R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}}),b^{\textit{BA}})>\textit{tf}(R^{A}_{i}+l^{A}_{i},\frac{p^{A}}{p^{B}}(R^{A}_{i}+l^{A}_{i}),b^{\textit{BA}})\,\,.$ (134)
From the definition of $i^{\ast}=i_{\min}(\textbf{R})$ (Equation 7.2.4), we have $R^{A}_{i^{\ast}}\leq R^{A}_{i}$. Combining with Equation 133, we have
$\frac{R^{A}_{i^{\ast}}}{\hat{l}^{A}_{i^{\ast}}+R^{A}_{i^{\ast}}}\leq\frac{R^{A}_{i}}{l^{A}_{i}+R^{A}_{i}}\,\,.$ (135)
Therefore, the liquidity provider has more share in $\textit{pool}_{i}$ in $\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ than any smallest pool in $\textbf{R}+a_{lp}$:
$\frac{\hat{l}^{A}_{i^{\ast}}}{\hat{l}^{A}_{i^{\ast}}+R^{A}_{i^{\ast}}}=1-\frac{R^{A}_{i^{\ast}}}{\hat{l}^{A}_{i^{\ast}}+R^{A}_{i^{\ast}}}\geq 1-\frac{R^{A}_{i}}{l^{A}_{i}+R^{A}_{i}}=\frac{l^{A}_{i}}{l^{A}_{i}+R^{A}_{i}}\,\,.$ (136)
Combining Equations 134 and 136, we have
$p^{B}\times\textit{tf}(R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}(R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}}),b^{\textit{BA}})\times\frac{\hat{l}^{A}_{i^{\ast}}}{\hat{l}^{A}_{i^{\ast}}+R^{A}_{i^{\ast}}}\geq\\ p^{B}\times\textit{tf}(R^{A}_{i}+l^{A}_{i},\frac{p^{A}}{p^{B}}(R^{A}_{i}+l^{A}_{i}),b^{\textit{BA}})\times\frac{l^{A}_{i}}{l^{A}_{i}+R^{A}_{i}}\,\,.$ (137)
Then, the revenue of the liquidity provider with the fill-up action is no less than any other action given the trader’s best response $\tau^{BA}$:
$U_{lp}(\textbf{R},a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}),a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}}))$
[…]
Therefore, from Equations 127 and 128, the revenue of the liquidity provider with the fill-up action is no less than any other action given the trader’s best response $\tau^{BA}$:
$U_{lp}(\textbf{R},a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}),\tau^{BA},\tau^{AB},l^{A},l^{B})\geq\\ U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$ (138)
After analyzing actions, we consider the utility of the liquidity provider with strategy $\tau_{lp}$ and any mixed strategy $\pi_{lp}$. From the definition of the utility of the liquidity provider under $\pi_{lp}$ (Equation 19), we have
$U_{lp}(\textbf{R},\tau_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})$
$=$ $\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\tau_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\times U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})$
$=$ $U_{lp}(\textbf{R},a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}),a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}})),$ (139)
and
$U_{lp}(\textbf{R},\pi_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})$
$=$ $\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\times U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})$
[…]
$=$ $U_{lp}(\textbf{R},a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B}),a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}}))$
×$\times\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\tau_{lp}(\textbf{R},l^{A},l^{B},a_{lp})$ (140)
$=U_{lp}(\textbf{R},\tau_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$ (141)
Therefore, the liquidity provider strategy $\tau_{lp}$ is the best response to the trader strategy $\tau^{BA}$:
$U_{lp}(\textbf{R},\tau_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\geq U_{lp}(\textbf{R},\pi_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\,\,.$ (142)
Since $\tau^{BA}$ is also a best response to the trader strategy $\tau_{lp}$, the liquidity provider strategy $\tau_{lp}$ and the trader strategy $\tau^{BA}$ are an SPNE. ∎
Appendix H Evaluation of Increasing Parallel Component
Figure 9: Maximal throughput as a function of the number of shards in SAMM / heavier SAMM.
We posit that enhancing the performance of SAMM necessitates mitigating the serial bottlenecks within the platform. While modifications to the core architecture of Sui are beyond the scope of this study, we enhance the parallelizable aspects of SAMM by introducing superfluous operations into each trade. This methodology follows the experimental setup described in Section 8.
Figure 9 presents the maximum throughput results for standard SAMM transactions as previously discussed in Section 8, alongside results from transactions with added operations with three repetitions. Although the inclusion of additional operations reduces overall performance due to increased overhead, it significantly enhances the parallel component, with $P=0.9$ ($R^{2}=0.968$). Remarkably, the throughput with $32$ shards is ten times that of a single shard, a substantial improvement over the ratios observed in Section 8.
Consequently, addressing the serial bottlenecks within blockchain platforms is vital for future improvements in SAMM performance.
Appendix I Parameter Selection in Simulation
We select different values of $c$ for SAMM, namely $0.003$, $0.005$ and $0.01$. We set $r_{\max}=5\times r_{\min}$ We choose $r_{\max},r_{\min}$ and $\beta_{1}$ (see Section 5.3) to minimize the maximal trading fee ratio. This is to limit the maximal cost for traders. Hence, we set $r_{\max}=5\times r_{\min}$ to minimize them at the same time. Together with the restrictions of satisfying $c$-smaller-better and $c$-larger-better (Corollary 5.3), the optimization problem is
$\underset{r_{\max},r_{\min},\beta_{1}}{\min}$ $r_{\max}$
subject to $r_{\max}=5\times r_{\min}\,\,,$
$c\leq 1-(-\beta_{1})^{-\frac{1}{3}}\,\,,$
$c\leq\frac{r_{\max}-r_{\min}}{-\beta_{1}}\,\,.$
Hence we get $r_{\max},r_{\min}$ and $\beta_{1}$ given $c$.